### v2

# Fully Energy-Efficient Randomized Backoff: Slow Feedback Loops Yield Fast Contention Resolution

Abstract.
Contention resolution addresses the problem of coordinating access to a shared communication channel. Time is discretized into synchronized slots, and a packet transmission can be made in any slot. A packet is successfully sent if no other packet is also transmitted during that slot. If two or more packets are sent in the same slot, then these packets collide and fail. Listening on the channel during a slot provides ternary feedback, indicating whether that slot had (0) silence, (1) a successful transmission, or (2+) noise. No other feedback or exchange of information is available to packets. Packets are (adversarially) injected into the system over time. A packet departs the system once it is successfully sent. The goal is to send all packets while optimizing throughput, which is roughly the fraction of successful slots.
Most prior contention resolution algorithms with constant throughput require a short feedback loop, in the sense that a packet’s sending probability in slot $t+1$ is fully determined by its internal state at slot $t$ and the channel feedback at slot $t$. This paper answers the question of whether these short feedback loops are necessary; that is, how often must listening and updating occur in order to achieve constant throughput? We can restate this question in terms of energy efficiency: given that both listening and sending consume significant energy, is it possible to have a contention-resolution algorithm with ternary feedback that is efficient for both operations?
A shared communication channel can also suffer random or adversarial noise, which causes any listener to hear noise, even when no packets are actually sent. Such noise arises due to hardware/software failures or malicious interference (all modeled as “jamming”), which can have a ruinous effect on the throughput and energy efficiency. How does noise affect our goal of long feedback loops/energy efficiency?
Tying these questions together, we ask: what does a contention-resolution algorithm have to sacrifice to reduce channel accesses? Must we give up on constant throughput? What about robustness to noise? Here, we show that we need not concede anything by presenting an algorithm with the following guarantees. Suppose there are $N$ packets arriving over time and $\mathcal{J}$ jammed slots, where the input is determined by an adaptive adversary. With high probability in $N+\mathcal{J}$, our algorithm guarantees $\Theta(1)$ throughput and $\operatorname{polylog}(N+\mathcal{J})$ channel accesses (sends or listens) per packet. We also have analogous guarantees when the input stream is infinite.
1. Introduction
Since the 1970s, randomized backoff protocols (Goldberg-notes-2000, ; HastadLeRo96, ; MetcalfeBo76, ; AbramsonKu73, ; 802.11-standard, ), such as binary exponential backoff (MetcalfeBo76, ), have been used for managing contention on a shared communication channel. Originally used in the ALOHA system (ALOHAnet, ; binder:aloha, ) and Ethernet (MetcalfeBo76, ; stuck1984introduction, ), randomized backoff plays an important role in a wide range of applications, including WiFi (802.11-standard, ), wireless sensor networks (khanafer2013survey, ), cellular systems  (cellular-standard, ), transactional memory (scherer2005advanced, ; HerlihyMo93, ), and congestion control (wierman2003unified, ). The salient feature of the communication channel is that it supports only one message transmission at a time: if more than one message is sent simultaneously, there is a collision resulting in indecipherable noise (Goldberg-notes-2000, ; kurose:computer, ; xiao:performance, ; DBLP:journals/jacm/GoldbergMPS00, ; MetcalfeBo76, ).
This contention-resolution problem is formalized as follows (hastad:analysis, ; goldberg2004bound, ; kelly1987number, ; aldous:ultimate, ; GoodmanGrMaMa88, ; mosely1985class, ; RaghavanU98, ; chlebus2009maximum, ; BenderGiKu22, ; zhou2019singletons, ; BenderFiGi19, ; FinemanNW16:multiple, ; fineman:contention, ; DeMarcoSt17, ; BenderFiGi16, ; FernandezAntaMoMu13, ; BenderFaHe05, ; deMarco:fast, ; DEMARCO20171, ; ChangJiPe19, ; BenderKoKu20, ; BenderKoPe16, ). There are $N$ packets arriving over time, and each packet needs to be successfully transmitted on the channel. [^1] For ease of exposition, we slightly abuse terminology and have the packets themselves taking action (e.g., sending themselves on the channel, listening on the channel), rather than introducing “agents”/“devices”/“senders”, where each one arrives on the scene with a packet to transmit. Time is divided into synchronized slots, each of which is sized to fit a single packet. To be successfully sent, the packet requires exclusive access to the channel (ChlamtacKu85, ; Goldberg-notes-2000, ; xiao:performance, ; DBLP:journals/jacm/GoldbergMPS00, ; hastad:analysis, ; MetcalfeBo76, ; AbramsonKu73, ); that is, the packet must be the only one transmitted during that slot. Otherwise, if two or more packets are transmitted in the same slot, the result is a collision, where none of the transmitted packets succeed. A packet departs the system once it is successfully sent. The objective is to send all packets while optimizing the throughput of the channel, which is roughly the fraction of successful slots.
The most famous contention-resolution protocol is binary exponential backoff (MetcalfeBo76, ) (see (kelly:decentralized, ; kelly1987number, ; rosenkrantz:some, ; aldous:ultimate, ; GoodmanGrMaMa88, ; GoodmanGrMaMa88, ; Al-Ammal2000, ; al2001improved, ; song:stability, )). Informally, under binary exponential backoff, a packet that has been in the system for $t$ time slots is sent with probability $\Theta(1/t)$.
Feedback loops: short versus long (versus no feedback).
An elegant, but ultimately problematic, feature of classical binary exponential backoff is that it is oblivious—packets remaining in the system do not use channel feedback to adjust their behavior: a packet with age $t$ sends with probability $\Theta(1/t)$ until it succeeds, regardless of channel history. The unfortunate result is that with adversarial packet arrivals, [^2] There is also much work analyzing the stochastic arrival rates under which binary exponential backoff is stable or unstable (HastadLeRo87, ; aldous:ultimate, ; GoodmanGrMaMa88, ); see also (MetcalfeBo76, ; Goldberg-notes-2000, ; hastad:analysis, ; goldberg2004bound, ; shoch1980measured, ; abramson1977throughput, ; kelly1987number, ; aldous:ultimate, ; GoodmanGrMaMa88, ; mosely1985class, ; RaghavanU98, ; RaghavanU95, ; GOLDBERG1999232, ). binary exponential backoff supports only a subconstant throughput—specifically, $O(1/\log{N})$; in fact, even for the batch case where all $N$ packets arrive at the same time, the throughput of binary exponential backoff is only $O(1/\log N)$ (BenderFaHe05, ).
In contrast, contention-resolution protocols can exploit frequent channel feedback to achieve $\Theta(1)$ throughput under adversarial arrivals (BenderKoKu20, ; BenderFiGi19, ; BenderFiGi16, ; BenderKoPe18, ; BenderKoPe16, ; ChangJiPe19, ; awerbuch:jamming, ). These protocols are not oblivious—packets listen on the channel and adjust their sending probabilities up or down based on this feedback.
The primary model for feedback on the channel is the ternary-feedback model (richa:jamming2, ; richa:jamming3, ; richa:jamming4, ; BenderFiGi19, ; DBLP:conf/spaa/AgrawalBFGY20, ; BenderFiGi16, ; BenderKoPe18, ; gilbert:near, ; DBLP:journals/dc/KingPSY18, ; chlebus2009maximum, ; chlebus2007stability, ; ChlebusKoRo06, ; fineman:contention2, ; anantharamu2019packet, ; greenberg1985lower, ). In this model, a packet can listen on the channel in each slot and learn whether that slot is (0) empty, if no packets send, (1) successful, if exactly one packet sends, or (2+) noisy, if two or more packets send. Based on this ternary feedback, the packet can decide when to attempt to send next. Once a packet succeeds, it immediately departs the system.
Most constant-throughput algorithms with ternary-feedback (e.g., (BenderKoKu20, ; BenderFiGi19, ; BenderFiGi16, ; BenderKoPe16, ; ChangJiPe19, ; DeMarcoSt17, ; awerbuch:jamming, ; richa:jamming2, ; richa:competitive-j, ; richa:jamming4, ; richa:jamming3, ; richa:efficient-j, ; ogierman:competitive, )) listen on the channel in every slot (or every constant number of slots). That is, these algorithms have a short feedback loop: a packet’s sending probability in slot $t+1$ is fully determined by its internal state at slot $t$ and the channel feedback at slot $t$. For example, the algorithm by Chang, Jin, and Pettie (ChangJiPe19, ), listens in every slot $t$ and multiplicatively updates the sending probability in slot $t+1$ based on whether it heard silence, a successful transmission, or a collision in slot $t$.
A key question is as follows: how frequently does the packet need to listen on the channel and update its behavior in order to achieve constant throughput? In every slot? In a vanishingly small fraction of slots? How short a feedback loop is necessary for good throughput?
As an analogy, one cannot navigate a ship without a control feedback loop: monitoring the surroundings and correcting course. One option is to constantly monitor and continuously update the heading to avoid obstacles. But the analogous question is whether one can still safely navigate with only a vanishingly small amount of course correction. In contention resolution, the monitoring is via the channel sensing and the course being corrected is the transmission probabilities.
Robustness and noise.
Finally, in addition to the factors discussed above, much of the recent work on contention resolution (anantharamu2019packet, ; anantharamu2011medium, ; Chlebus:2016:SWM:2882263.2882514, ; DBLP:conf/spaa/AgrawalBFGY20, ; ChenJiZh21, ; jiang:robust, ; awerbuch:jamming, ; richa:jamming2, ; richa:competitive-j, ; richa:jamming4, ; richa:jamming3, ; richa:efficient-j, ; ogierman:competitive, ; BenderFiGi19, ; BenderFiGi16, ; ChangJiPe19, ) has sought to address an additional factor, which is that the real world is often noisy. Sometimes interference prevents transmissions in a slot and listeners hear noise even if nobody actually sends. Noisy channels arise due to hardware/software failures, co-located devices, or malicious jamming (chiwewe2015using, ; liu2017interference, ; pelechrinis:denial, ; mpitziopoulos:survey, ; xu:feasibility, ). Regardless of the source of the noise in a slot, we may think of these slots as being “jammed” by an adversary. Jamming has evolved from a mostly-theoretical risk into a credible threat to systems over the past decade, with several publicized examples (nicholl, ; FCC:humphreys, ; marriot, ; priest, ).
After a long line of work, there are now many contention-resolution protocols that achieve constant throughput in the presence of noise (e.g. (awerbuch:jamming, ; richa:jamming2, ; richa:competitive-j, ; richa:jamming4, ; richa:jamming3, ; richa:efficient-j, ; ogierman:competitive, ; BenderFiGi19, ; BenderFiGi16, ; ChangJiPe19, )); however, again, these protocols listen to the channel in every slot and update their sending probabilities accordingly. Our goal is to eliminate short feedback loops not just in the classical model, but also with jamming.
Minimizing channel accesses = energy efficiency.
Up until now, we have discussed whether one can minimize listening and thus avoid providing immediate feedback for a contention-resolution algorithm. Another way of viewing this problem is through the lens of energy efficiency. Each channel access—whether for sending or listening—consumes energy. Most work on contention resolution is sending efficient, but is not listening efficient (e.g.,  (awerbuch:jamming, ; richa:jamming2, ; richa:competitive-j, ; richa:jamming3, ; richa:efficient-j, ; richa:jamming4, ; BenderFiGi19, ; BenderKoKu20, ; DeMarcoSt17, ; jiang:robust, ; de2022contention, )). That is, most protocols optimize how frequently a packet sends on the channel, but allow a packet to listen on the channel in every slot “for free”.
In fact, both sending and listening are expensive operations (e.g., (feeney2001investigating, ; polastre:telos, ; yang2017narrowband, ; lauridsen2018empirical, )), and minimizing energy usage by having devices sleep as much as possible has been a long-standing and popular strategy to maximize the network lifetime (e.g., (merlin2010duty, ; ye2006ultra, ; li2005energy, ; deng2005scheduling, )).333 Why bother optimizing listening in contention resolution protocols if packets must listen to receive messages (for other network applications)? Informally, there are two types of listening: listening to receive messages, and listening to execute a contention resolution protocol (which is what enables sending messages). In order to receive messages, packets do not need to listen in every time slot; there exist methods for optimizing this first type of listening. Although this is beyond the scope of this paper, as one example, in many wireless settings, a central base station can monitor the channel constantly and facilitate message exchange (feng2012survey, ). We call a protocol fully-energy efficient if it is both sending efficient and listening efficient. By definition, such a protocol cannot have short feedback loops, since it can access the channel only rarely.
Past work: Minimizing listening by allowing for explicit synchronization.
Currently, the only known path to full energy efficiency is via explicit synchronization. This means that the model is extended so that packets can send synchronization messages to each other whenever they broadcast (BenderKoPe18, ; DeMarcoSt17, ; marco:time, ). In (BenderKoPe18, ), these synchronization messages have size $\Theta(\log N)$ bits each, which means that an arbitrary polynomial amount of communication can be expressed in a slot. In (DeMarcoSt17, ; marco:time, ), the synchronization messages are smaller ($O(1)$ bits), but some packets are permitted to stick around as “Good Samaritans” in order to serve as long-term coordinators (that send many messages via many broadcasts).
Using $\Theta(\log N)$-bit synchronization messages, Bender, Kopelowitz, Pettie, and Young (BenderKoPe18, ) give an algorithm with $\Theta(1)$ throughput and expected $O(\log(\log^{*}N))$ channel accesses per packet. Using Good-Samaritan packets, De Marco and Stachowiak (DeMarcoSt17, ) and De Marco, Kowalski, and Stachowiak (marco:time, ) provide a constant-throughput algorithm that is sending efficient ($O(\log N)$ transmissions per packet) and conjecture that their techniques can be extended to achieve fully energy efficiency ($O(\operatorname{polylog}N)$ channel accesses per packet), with high probability and even without collision detection.
In all of these cases, even with the help of explicit synchronization, it remains open whether one can achieve such results in the presence of adversarial noise. Indeed, noise has the potential to be extra-problematic for energy-efficient algorithms since these algorithms listen to the channel less frequently and can potentially be thrown off by a small amount of well-placed noise.
This paper.
We show that it is indeed possible to achieve constant throughput with full energy efficiency (i.e., both listening efficiency and sending efficiency) while supporting robustness to adversarial noise. Moreover, these results hold in the standard ternary-feedback model, without requiring the addition of any sort of explicit synchronization.
Our algorithm belongs to a natural family of multiplicative-weight-update algorithms (e.g., (ChangJiPe19, ; awerbuch:jamming, ; richa:jamming2, ; richa:jamming3, ; ogierman:competitive, ; richa:jamming4, )). When a packet hears silence, it multiplicatively increases both its listening and sending probabilities. Conversely, when a packet hears noise, it multiplicatively decreases these probabilities. There are no control messages, no Good Samaritan packets, and no leaders elected.
What makes our algorithm/analysis interesting is that we are able to support a multiplicative-weight-update framework while having each packet ‘cover its eyes’ almost all of the time. This is in stark contrast to prior constant-throughput algorithms, which adjust sending probabilities in every slot. Because each packet listens to so few slots, the different packets that are in the system at the same time may end up with very different perspectives on the world from each other. In order to analyze the ‘herd behavior’ of the packets in this potentially chaotic setting, substantially new techniques end up being required. These techniques are also what allow us to handle the additional chaos that adversarial jamming (i.e., noise) further adds to the system.
1.1. Model
A finite or infinite stream of indistinguishable packets arrives over time; the number of arrivals is unknown to the algorithm. Each packet needs to be sent on the multiple-access channel. Time is divided into synchronized slots, each of which is sufficiently large to send a single packet. An adversary (specified below) controls precisely how many packets are injected into the system in each slot. When a packet successfully transmits, it leaves the system. There is no universal numbering scheme for the slots; that is, there is no global clock from which a newly injected packet could infer the lifetime of the system or slot parity. Additionally, the packets do not receive any additional information about how many packets have arrived or will arrive. Instead, packets only receive information through the ternary feedback model.
We now describe the ternary feedback model (richa:jamming2, ; richa:jamming3, ; richa:jamming4, ; BenderFiGi19, ; DBLP:conf/spaa/AgrawalBFGY20, ; BenderFiGi16, ; BenderKoPe18, ; gilbert:near, ; DBLP:journals/dc/KingPSY18, ; chlebus2009maximum, ; chlebus2007stability, ; ChlebusKoRo06, ; fineman:contention2, ; anantharamu2019packet, ; greenberg1985lower, ; willard:loglog, ). Initially, we will define the model without jamming. In each time slot, each packet in the system can take one of three actions: (i) sleep, (ii) send, or (iii) listen to the channel. Packets that take actions (ii) or (iii) are said to access the channel. If no packets choose to send during a slot, then that (non-jammed) slot is empty/silent; if exactly one packet sends, then that (non-jammed) slot is full and successful; if two or more packets send, then that slot is full and noisy. A packet that listens during a slot (action iii) learns whether the slot was (0) empty, (1) successful, or (2+) noisy. A packet that sleeps during a slot (action i) learns nothing about the state of the slot. A packet that sends (action ii), either transmits successfully and leaves the system, or collides and remains in the system. [^4] For ease of presentation, in our algorithms, we say that a packet can listen and send simultaneously, but any packet that is sending actually does not need to listen to determine the state of the channel. If the packet is still in the system after sending in slot $t$, then slot $t$ was noisy.
We now add jamming to the picture; an adversary determines which slots are jammed. To jam a particular slot, the adversary broadcasts noise into that slot. All jammed slots are thus full and noisy. A packet that listens in a jammed slot hears that the slot was noisy, but does not know whether that noise came from jamming or was merely a collision between two or more packets. A packet that sends during a jammed slot collides and thus remains in the system.
An adversary determines, for each slot $t$, how many packets to inject in slot $t$ and whether to jam in that slot. This paper considers an adaptive adversary, which bases its decision on the entire state of the system so far, i.e., up to the end of slot $t-1$, but not the outcomes of future coin tosses. Thus, if at slot $t$, a packet $p$ decides whether to send based on a coin flip, the adaptive adversary does not get to see that coin flip until after slot $t$.
The basic metric: (overall) throughput. The main objective of contention resolution is to optimize throughput, defined next. A slot is active if at least one packet is in the system during that slot; inactive slots can be completely ignored in our analysis. Without loss of generality, assume throughout that the first slot is active. Without jamming, the throughput at time $t$ is defined as $T_{t}/S_{t}$, where $T_{t}$ is the number of successful transmissions during slots $1,2,\ldots,t$, and $S_{t}$ is the number of active slots in the same time interval. [^5] By assumption that the first slot is active, we have $S_{t}\geq 1$ and hence throughput is always well defined. On a finite input, the (overall) throughput is defined with respect to the final active slot $t$; at this point, the overall throughput is $N/S$, where $N=T_{t}$ is the total number of packets and $S=S_{t}$ is the total number of active slots. Overall throughput is not well-defined on an infinite execution.
A deficiency of the throughput metric (when defined naïvely) is that even if an algorithm guarantees $\Theta(1)$ overall throughput, it is not possible to achieve $\Theta(1)$ throughput uniformly across time. For example, if there is a burst of $N$ packets at time 00, and $N$ is unknown to the algorithm, then the throughput will be 00 for a superconstant number of slots (e.g., see (willard:loglog, ; ChangKoPe19, )), and this is provably unavoidable regardless of the backoff strategy being used.
Thus (overall) throughput is only meaningful at the end of the execution, or at points in time where there are no packets in the system.
A stronger metric: implicit throughput (BenderKoKu20, ). In order to support a meaningful notion of throughput, even at intermediate points in time, Bender, Kopelowitz, Kuszmaul, and Pettie (BenderKoKu20, ) propose a refined definition that they term “implicit throughput” (BenderKoKu20, ). The implicit throughput at time $t$ is defined as $N_{t}/S_{t}$, where $N_{t}$ is the total number of packets that arrive at or before time $t$ and $S_{t}$ is the total number of active slots so far. [^6] Notice that while $N_{t}$ depends on the adversary, the number of active slots depends on the contention resolution algorithm, whose goal it is to make slots inactive by completing packets.
One perspective on implicit throughput is that it is an analytical tool. Indeed, whenever we reach a point in time where overall throughput is meaningful (i.e., there are no packets left in the system), the two metrics become provably equal. This includes both at the end of any finite execution or during quiet periods of infinite executions. Another perspective on implicit throughput is that it is a stronger metric that offers a meaningful guarantee even at intermediate points in time: what constant implicit throughput means is that the number of active slots used so far should never be asymptotically larger than the number of packets that have arrived.
Throughout the rest of the paper, we will focus exclusively on implicit throughput. We should emphasize, however, that this only makes our results stronger—the results also imply the standard constant-throughput guarantees that one would normally strive for.
Extending to adversarial jamming. We next extend the definitions of throughput and implicit throughput for the case of jamming following (BenderFiGi16, ; BenderFiGi19, ). An algorithm wastes a slot if that slot has silence or a collision, and throughput measures the fraction of slots that the algorithm could have used but instead wasted. Let $\mathcal{J}_{t}$ denote the number of jammed slots through slot $t$. Then, the throughput of an execution ending at time $t$ is defined to be $(T_{t}+\mathcal{J}_{t})/S_{t}$, and the implicit throughput at slot $t$ is defined to be $(N_{t}+\mathcal{J}_{t})/S_{t}$.
Some useful properties of implicit throughput, and applications to adversarial queuing theory.
We conclude the section by summarizing several useful properties of implicit throughput.
Observation 1 ( (BenderKoKu20, )).
Consider any inactive slot $t$, i.e., where there are no active packets in the system. Then the implicit throughput and throughput are the same at slot $t$.
Observation 2 ( (BenderKoKu20, )).
Let $\delta$ be any lower bound on the implicit throughput of an algorithm; that is, suppose the algorithm achieves implicit throughput of at least $\delta$ at all times. Let $N_{t}$ and $S_{t}$ denote the total number of packet arrivals and active slots, respectively, at or before time $t$. Then $S_{t}\leq N_{t}/\delta$. Consequently
  • Overall throughput. Suppose that there are a total of $N\geq 1$ packet arrivals. Then the total number of active slots is at most $N/\delta$, and hence the overall throughput is at least $\delta$.
  • Backlog reprieve. If $N_{t}<\delta t$, then there exists an inactive slot $t^{\prime}\leq t$. Thus, all packets that arrived before $t^{\prime}$ have completed before $t^{\prime}$, and hence the throughput at time $t^{\prime}$ is at least $\delta$.
Finally, we observe that there are several natural settings in which implicit-throughput guarantees directly imply strong guarantees on packet backlog, even for infinite input sequences. Suppose, in particular, that packets arrive according to adversarial queuing theory, which parameterizes the “burstiness” of packet-arrival in infinite streams. In adversarial queuing theory (BenderFaHe05, ; ChlebusKoRo12, ; BKS14, ; ChlebusKoRo06, ; BorodinKlRaSu01, ; BorodinKlRa96, ), the adversary is restricted from injecting too many packets/jammed slots over a set of consecutive slots of length $S$, where $S$ is a parameter of the model that we refer to as the granularity. For granularity $S$, the number of packet arrivals plus the number of jammed slots is limited to $\lambda S$, where the arrival rate $\lambda$ is a constant less than one. On the other hand, how the packet arrivals are distributed within each $S$-sized window is adversarial: no restrictions are placed on how the (at most) $\lambda S$ packets and jammed slots are distributed. By showing that the implicit throughput of all active slots is $\Omega(1)$, we obtain as a corollary a strong bound of $O(S)$ on the number of packets backlogged in the system at all times, as long as $\lambda$ is a sufficiently small constant.
1.2. Main results
We present an algorithm that with high-probability guarantees $\Omega(1)$ implicit throughput in all active slots and full energy efficiency. Thus, we resolve two open questions in contention resolution: First, we show that full-energy efficiency is feasible, even with only ternary feedback. Second, we show that these guarantees are achievable, even in the presence of adversarial jamming.
Our algorithm relies on the natural multiplicative-weight approach to backoff—with a careful choice of probabilities and updates rules. In contrast to some previous approaches, we do not rely on packet batching (to turn the online problem into a series of batch problems), leader election, busy tones, population estimation, dividing the channel into simulated subchannels (e.g., odd and even slots), or other approaches seen in many modern algorithms; alas, these seem problematic in our setting (and potentially complicated to implement). Given the simplicity of our algorithm, the technical innovation lies in choosing the update parameters, analyzing the underlying combinatorial process, and proving that it is fast, robust, and fully energy efficient.
Our algorithm Low-Sensing Backoff guarantees the main theorems below. [^7] An event occurs with high probability (w.h.p.) in $x$ if for any fixed constant $c\geq 1$, the probability of the event is at least $1-x^{-c}$.
Theorem 3 (Implicit throughput, infinite packet streams).
At the $t$-th active slot, the implicit throughput is $\Omega(1)$ w.h.p. in $t$.
Corollary 4 (Throughput, finite packet streams).
Consider an input stream of $N$ packets with $\mathcal{J}$ jammed slots. The throughput for the execution is $\Theta(1)$ w.h.p. in $N+\mathcal{J}$.
Corollary 5 (Bounded backlog for adversarial-queuing arrivals).
Consider adversarial-queuing-theory arrivals with a sufficiently small constant arrival-rate $\lambda$ and granularity $S$; i.e., in any interval of length $S$, the total number of packet arrivals and jammed slots is at most $\lambda S$. Then, for any given slot $t$, the number of packets currently in the system is at most $O(S)$ w.h.p. in $S$.
Theorem 6 (Energy, finite executions).
Consider a finite execution with $N$ total packet arrivals and $\mathcal{J}$ total jammed slots against an adaptive adversary. Any given packet accesses the channel $O(\operatorname{polylog}(N+\mathcal{J}))$ times w.h.p. in $N+\mathcal{J}$.
Theorem 7 (Energy, adversarial queuing).
Consider a (finite or infinite) packet stream with adversarial-queuing arrivals with granularity $S$ and where the arrival rate is a sufficiently small constant. Then any given packet accesses the channel $O(\operatorname{polylog}(S))$ times w.h.p. in $S$ against an adaptive adversary.
Theorem 8 (Energy, infinite executions).
Consider an infinite packet stream, and let $N_{t}$ and $\mathcal{J}_{t}$ denote the number of arrivals and jammed slots, respectively, up until time $t$ against an adaptive adversary. Then any given packet accesses the channel $O(\operatorname{polylog}(N_{t}+\mathcal{J}_{t}))$ times before time $t$, w.h.p. in $N_{t}+\mathcal{J}_{t}$.
Theorem 3 is proved in Section 5.5 where it appears as Corollary 11. Corollary 5 corresponds to Corollary 14. The remaining theorems are proved in Section 5.6. In particular, Theorem 6 corresponds to Theorem 3.4, and Theorem 7 corresponds to Theorem 16. Finally, Theorem 8 is included in Theorem 3.4.
1.3. Extensions to Reactive Adversary
A reactive adversary (richa:efficient-j, ; wilhelm2011short, ; DBLP:journals/dc/KingPSY18, ) has an instantaneous reaction time; that is, this adversary listens to the channel and can decide whether to jam and/or inject new packet(s) in slot $t$ based on what it hears in slot $t$ itself. In contrast, the standard adaptive adversary would not know whether any packet chooses to send in slot $t$ until slot $t+1$. This allows a reactive adversary to cheaply prevent any particular packet $p$ from transmitting successfully by jamming only those slots where $p$ makes transmission attempts. Consequently, against a reactive adversary, the total number of channel accesses required is at least linear in the amount of jamming. In the case of exponential backoff, the situation is more dire: for any $T$ a reactive adversary can also drive the throughput down to $O(1/T)$ by jamming a single packet a mere $\Theta(\log T)$ times.
Note that reactivity and adaptivity are somewhat orthogonal. Reactivity addresses how quickly the adversary can react to the visible state on the channel—importantly, only sending is revealed in this way, since it is visible. In contrast, the adaptive adversary knows all of the internal state and random choices of packets up to the previous slot, and in particular this adversary also knows if packets choose to listen.
In addition to our main results on purely adaptive adversaries, this paper also addresses an adversary that is both adaptive and reactive. It turns out that the reactive adversary does not impact the implicit-throughput bounds for our algorithm (our analysis applies whether or not the adversary can see the activity on the channel at the current time). Reactivity thus only impacts the number of channel accesses. Roughly speaking, the theorem states that the reactive adversary has nontrivial impact on the worst-case number of channel accesses and thus energy, which is to be expected, but it does not have significant impact on the average.
Theorem 9 (Energy, reactive adversary).
The following apply to a reactive and adaptive adversary.
  1. (1)
    Finite streams. Consider a finite execution with $N$ total packet arrivals and $\mathcal{J}$ total jammed slots. Any given packet accesses the channel $O((\mathcal{J}+1)\operatorname{polylog}(N))$ times w.h.p. in $N+\mathcal{J}$. Moreover, the average number of channel accesses is only $O((\mathcal{J}/N+1)\operatorname{polylog}(N+\mathcal{J}))$ times w.h.p. in $N+\mathcal{J}$.
  2. (2)
    Adversarial queuing. Consider a (finite or infinite) packet stream with adversarial-queuing arrivals with granularity $S$ and where the arrival rate is a sufficiently small constant. Then any given packet accesses the channel at most $O(S)$ times, w.h.p. in $S$. In addition, the average number of accesses per slot is $O(\operatorname{polylog}(S))$, w.h.p. in $S$.
  3. (3)
    Infinite executions. Consider an infinite packet stream, and let $N_{t}$ and $\mathcal{J}_{t}$ denote the number of arrivals and jammed slots, respectively, up until time $t$. Then any given packet accesses the channel $O((\mathcal{J}_{t}+1)\operatorname{polylog}(N_{t}+\mathcal{J}_{t}))$ times before time $t$, w.h.p. in $N_{t}+\mathcal{J}_{t}$. Moreover, the average number of channel accesses is $O((\mathcal{J}_{t}/N_{t}+1)\operatorname{polylog}(N_{t}+\mathcal{J}_{t}))$.
The items of Theorem 9 are each proved separately as Theorems 15, 17, and 3.4 in Section 5.6.
2. Low-Sensing Backoff Algorithm
This section presents the Low-Sensing Backoff algorithm; see Figure 1.
For ease of presentation, we describe our algorithm as listening whenever it sends. However, the packet need not actually do both; observe that any packet that is sending does not need to listen to determine the state of the channel, since if the packet is still in the system after sending in slot $t$, then slot $t$ was noisy.
Low-Sensing Backoff for packet $u$ Key Variables: $w_{u}(t)$: window size of $u$ in slot $t$. If $u$ is injected at time slot $t$, then $w_{u}(t)=w_{\rm min}$. $c$: a sufficiently large positive constant. In every slot $t$, packet $u$ executes the following four steps with probability $\frac{c\ln^{3}\left(w_{u}(t)\right)}{w_{u}(t)}$: Listen Send with probability $\frac{1}{c\ln^{3}(w_{u}(t))}$ If $u$ heard a silent slot, then $w_{u}(t+1)\leftarrow\max{\left\{\mbox{\small$\displaystyle\frac{w_{u}(t)}{1+1/(c\ln(w_{u}(t)))}$},\,w_{\rm min}\right\}}$ If $u$ heard a noisy slot, then $w_{u}(t+1)\leftarrow w_{u}(t)\cdot\left(\mbox{\small$\displaystyle 1+\frac{1}{c\ln(w_{u}(t))}$}\right)$
Figure 1. Low-Sensing Backoff algorithm.
The probabilities for sending and listening in Low-Sensing Backoff are determined by a single parameter, which we call packet $u$’s window size. Let $w_{u}(t)$ denote packet $u$’s window size at time slot $t$. When $u$ is injected into the system, its window size is set to the minimum allowed value: $w_{\rm min}$ $\geq 2$. Let $c$ be a sufficiently large positive constant. The sending and listening rules are as follows. First, packet $u$ listens with probability $c\ln^{3}\left(w_{u}(t)\right)/w_{u}(t)$. Then, conditioned on listening, $u$ sends with probability $1/(c\ln^{3}(w_{u}(t)))$.
A packet $u$ only has the option to change its window size when it accesses the channel. Specifically, if at time $t$, packet $u$ listens to the channel and learns that the slot $t$ is busy, then the window size increases (or backs off) by a backoff factor of $1+1/(c\ln(w_{u}(t)))$; that is, $w_{u}(t+1)\leftarrow w_{u}(t)(1+1/(c\ln(w_{u}(t))))$. Similarly, if at time $t$, packet $u$ accesses the channel and learns that the slot $t$ is empty, then the window size shrinks (or backs on) by a backon factor of $1+1/c\ln(w_{u}(t))$, or until it gets back down to $w_{\rm min}$, that is, $w_{u}(t+1)\leftarrow\max\left\{w_{u}(t)/(1+1/(c\ln w_{u}(t))),\,w_{\rm min}\right\}$.
3. Technical Overview
This section gives a technical overview. In Section 3.1, we introduce the notion of contention. In Section 3.2, we introduce our potential function $\Phi(t)$. Section 3.3 gives the main structure of our analysis in terms of intervals. Finally, Section 3.4 provides a synopsis of the main analytical results achieved and how they are deployed to make progress towards our main results (stated in Section 1.2). That is, we describe the main point of each of the technical sections, namely Sections 5.15.6.
3.1. Contention
For any slot $t$, we define the contention $C(t)$ $=\sum_{u}1/w_{u}(t)$ to be the sum of the sending probabilities in that slot, i.e., the expected number of packets that attempt to send during that slot. We say contention is high when $C(t)>C_{\rm high}$, where $C_{\rm high}$$>1$ is some fixed positive constant. Conversely, we say that contention is low when $C(t)<C_{\rm low}$, where we define $C_{\rm low}$ to be some fixed positive constant such that $C_{\rm low}\leq 1/w_{\rm min}$. Otherwise, if contention is in $[C_{\rm low},C_{\rm high}]$, then we say that contention is good.
3.2. Our potential function
Throughout the execution of Low-Sensing Backoff, we maintain a potential function $\Phi(t)$ that captures the state of the system at time $t$ and measures the progress toward delivering all packets. When a slot $t$ is inactive, $\Phi(t)=0$. We will see that packet arrivals increase $\Phi(t)$ by $\Theta(1)$ per newly arrived packet, that packets exiting the system successfully decrease $\Phi(t)$ by $\Theta(1)$ per packet, that a jammed slot increases $\Phi(t)$ by $O(1)$, and that on average each slot decreases the potential by $\Theta(1)$, ignoring newly arrived packets.
For any slot $t$, $N(t)$ is the number of packets in the system, $w_{u}(t)$ is $u$’s window size, $w_{\max}(t)$ is the largest window size over all packets, and $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are positive constants. Our potential function consists of three terms. Implicitly, the third term is 0 if there are no packets in the system (and thus $w_{\max}(t)=0$):
${\boldmath{\Phi(t)}}=\alpha_{1}N(t)\,+\,\alpha_{2}\sum_{u}\frac{1}{\ln(w_{u}(t))}\,+\,\alpha_{3}\frac{w_{\max}(t)}{\ln^{2}(w_{\max}(t))}\,.$
We abbreviate $\Phi(t)$ as:
$\Phi(t)=\alpha_{1}N(t)+\alpha_{2}H(t)+\alpha_{3}L(t),$
where $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ may be set so that $\Phi(t)$ will decrease as time progresses for all values of contention $C(t)$. The notation $H(t)$ and $L(t)$ is used to highlight that these terms capture the impact on $\Phi$ from high contention and low contention, respectively.
Why these terms?
There are three main features of the state of the system that are captured by the potential function: the number of packets, the contention, and the size of the windows. (Note that these are not independent, as larger windows correspond to lower contention.) Roughly speaking, when there are many active packets, potential should be high, and when there are no packets, the potential should be 0. The $N(t)$ term captures this idea directly by counting the number of packets.
The $H(t)$ term is chosen so that the expected change to $H(t)$ in a slot is proportional to the contention. When the contention is high (and the slot is most likely to have a collision), in expectation $H(t)$ decreases proportional to the contention (due to the update rule on noisy slots). On the other hand, when the contention is low (and the slot is most likely to have silence), $H(t)$ increases proportional to the contention. Overall, this is pretty great: when contention is high, $H(t)$ is likely to decrease by a lot. When contention is low, there is a small expected increase, but that increase is counterbalanced by the (small) expected number of packet successes reflected in $N(t)$. Choosing $\alpha_{1}>\alpha_{2}$ makes the net effect a decrease.
Finally, the $L(t)$ term allows us to cope with the situation that the contention is low but some packets in the system have large windows (e.g., there is a single packet with a very large window). As it is likely to take a long time for the packet to succeed, the potential should be high. $L(t)$ is roughly the expected time for a packet with window $w_{\max}(t)$ to decrease its window size to a constant if all slots are silent. The analysis then needs to show that any increases to $L(t)$ are counterbalanced by decreases in the other terms, ensured by $\alpha_{1}>\alpha_{2}>\alpha_{3}$.
Challenge with $L(t)$.
The $N(t)$ and $H(t)$ terms are well-behaved in the sense that they change on a per-slot basis, while the $L(t)$ term cannot. To see why, consider the case that several packets with window size $w_{\max}(t)$ remain. The $L(t)$ term only decreases after all of those packets have chosen to listen and observed silence. On any step that multiple such packets remain, it is extremely unlikely that all of the packets choose to listen. Thus, $L(t)$ does not decrease by a constant in expectation. Instead, we need a coarser granularity to understand the behavior of $L(t)$.
3.3. Analyzing intervals
Our analysis divides the execution into disjoint intervals of time. The first interval starts at the first step with an active packet. An interval starting at time $t$ has size $\tau=(1/c_{\tau})\max\{L(t),\sqrt{N(t)}\}$, where $c_{\tau}$ is a constant. If any active packets remain, the next interval starts immediately after the previous interval ends. (Otherwise, an interval begins the next time there is an active packet.)
A key technical theorem is the following. Let $\mathcal{A}$ and $\mathcal{J}$ denote the number of arrivals and jammed slots, respectively, in the size-$\tau$ interval. For $\mathcal{A}=\mathcal{J}=0$, the lemma states that the potential decreases by $\Omega(\tau)$ across the interval, with high probability in $\tau$, meaning a decrease of $\Omega(1)$ per slot. For general $\mathcal{A},\mathcal{J}$, the potential decreases by $\Omega(\tau)-O(\mathcal{A}+\mathcal{J})$.
{restatable*}
[Decrease in $\Phi(t)$ over interval $\mathcal{I}$ w.h.p. in $|\mathcal{I}|$]theoremphimain Consider an interval $\mathcal{I}$ starting at $t$ of length $|\mathcal{I}|=\tau=(1/c_{\tau})\cdot\max\big{\{}\frac{w_{\max}(t)}{\ln^{2}(w_{\max}(t))},N(t)^{1/2}\big{\}}$. Let $\mathcal{A}$ and $\mathcal{J}$ be the number of packet arrivals and jammed slots in $\mathcal{I}$. With high probability in $\tau$, $\Phi$ decreases over $\mathcal{I}$ by at least $\Omega(\tau)-O(\mathcal{A}+\mathcal{J})$. That is,
$\Pr\left[(\Phi(t^{\prime})-\Phi(t))\,\geq\,\Theta(\mathcal{A}+\mathcal{J})-\Omega(\tau)\right]\,\leq\,(1/\tau)^{\Theta(1)}\ .$
Our proof of Theorem 3.3 is broken into several lemmas according to the level of contention. Specifically, we have separate cases for high contention, good contention, and low contention. In each of the cases, absent arrivals and jamming, we argue that there is a net decrease in potential, with high probability, but the contributing term is different in each case. The interplay between $N(t)$ and $H(t)$ is tight enough that we analyze the net effect on the sum of these terms together, but we analyze $L(t)$ separately. A more detailed summary is provided next in Section 3.4.
A significant complication is that (1) the probability stated in Theorem 3.3 depends on the size of the interval, and (2) the interval sizes are determined adaptively by actions of the adversary. To analyze the full process, we model an execution as a specific biased random walk that we set up as a betting game (Section 5.5). The bounds provided by the betting game translate into high-probability bounds with respect to the total number of packets.
Throughout the paper, standard Chernoff bounds sometimes cannot be used for two reasons. First the adversary can adaptively influence the length of an interval. Moreover within each interval, the adversary can influence which slots are high, low, and good contention. To be able to analyze these slots separately, we must instead apply a generalization of Azuma’s inequality (see Theorems 4 and 5 in Section 5.1, taken from (KuszmaulQi21, )) that gives us Chernoff-like bounds but with adaptively chosen probability distributions.
Finally, good upper bounds for $\Phi(t)$ enable us to characterize the (implicit and standard) throughput and energy consumption of Low-Sensing Backoff in all its variety of settings (finite versus infinite executions, arbitrary infinite versus infinite with adversarial-queuing arrivals, adaptive adversaries that are reactive versus non-reactive). The most direct application of $\Phi(t)$ is to bound implicit throughput. $\Phi(t)$ also gives us an upper bound on the maximum window size $w_{\max}(t)$, specifically, $w_{\max}(t)=O\big{(}\Phi(t)\log^{2}(\Phi(t))\big{)}$, which we use to prove energy bounds in Section 5.6.
3.4. Proof organization
The main analysis in this paper, including all of the proofs, appears in Section 5 in our Appendices A and B. This section summarizes the proof structure, highlighing the key lemma statements.
Overview of Section 5.1.  Preliminaries.
This section lists several well-known inequalities (Fact 1) that are used throughout our analysis. We review bounds on the probability that a slot is noisy, empty, or contains a successful transmission as a function of contention (Lemmas 1, 2 and 3). The lemmas in this section allow us to immediately obtain constant bounds on the probabilities of empty slots, successful slots, and noisy slots in different contention regimes.
Theorems 4 and 5 provide upper and lower bounds for the sum of random variables, where the distribution of each subsequent random variable is determined by an adaptive adversary. This adversarial, multiplicative version of Azuma’s inequality is a powerful probabilistic tool from (KuszmaulQi21, ) that allows us to analyze the performance of our algorithm in situations where a simpler Chernoff-bound-style argument does not appear to work, given the adaptive nature of our adversary.
[…]
This section addresses the behavior of $N(t)$ and $H(t)$. When contention is high, we expect to see a decrease in $H(t)$, which should be large enough that its reduction outweighs any increase from $L(t)$, and thus $\Phi(t)$ decreases.
Lemma 3.4 shows how much $H(t)$ changes as a result of a specific packet listening during a slot $t$—that is, how much $H(t)$ increases when the slot is silent and decreases when the slot is noisy. Lemma 3.4 analyzes the change to $N(t)+H(t)$ due to the low and good contention slots in an arbitrary interval. We highlight that the adaptive adversary exerts some control over which slots have low and good contention, since it can inject packets and/or jam in slot $t+1$ based on the packets’ random choices in slot $t$. In particular, let $|\mathcal{G}|$ denote the number of good slots in the interval, then Lemma 3.4 shows the following. Over good-contention slots, $N(t)+H(t)$ decreases by $\Omega(|\mathcal{G}|)$, minus the number of packet injections, jammed slots, and a polylog term in the length of the interval, w.h.p. in the interval length. Lemma 3.4 also shows that over the low-contention slots, $N(t)+H(t)$ increases by at most the number of packet injections, jammed slots, and a polylog term in the length of the interval, again w.h.p. in the interval length.
{restatable*}
[Increase/decrease in $H(t)$ due to a silent/noisy slot]lemmaBterminc When packet $u$ listens to a silent slot $t$, $H(t)$ increases by $\Theta\big{(}\frac{1}{c\ln^{3}w_{u}}\big{)}$ due to packet $u$. When a packet $u$ listens to a noisy slot, $H(t)$ decreases by $\Theta\big{(}\frac{1}{c\ln^{3}w_{u}}\big{)}$ due to packet $u$.
{restatable*}
[Net delta in contribution of $N(t)$ and $H(t)$ to potential over low and good contention slots]lemmaABlowmid Let $\mathcal{I}$ be an arbitrary interval starting at time $t$ with length $|\mathcal{I}|=\tau$. Let $\mathcal{L}$ be the set of slots in $\mathcal{I}$ during which $C(t)\leq C_{\rm low}$. Let $\mathcal{G}$ be the set of slots in $\mathcal{I}$ during which $C(t)>C_{\rm low}$ and $C(t)\leq C_{\rm high}$. Let $\mathcal{A}_{\mathcal{L}}$ be the number of packet arrivals in time slots in $\mathcal{L}$. Let $\mathcal{A}_{\mathcal{G}}$ be the number of packet arrivals in time slots in $\mathcal{G}$. Let $\mathcal{J}_{\mathcal{L}}$ be the number of jammed slots in $\mathcal{L}$. Let $\mathcal{J}_{\mathcal{G}}$ be the number of jammed slots in $\mathcal{G}$. Define:
  • the net delta over $\mathcal{L}$ to be the sum of the changes in $N(t)$ and $H(t)$ during the slots in $\mathcal{L}$, i.e.,
    $\sum_{t^{\prime}\in\mathcal{L}}\Big{(}\alpha_{1}\big{(}N(t^{\prime}+1)-N(t^{\prime})\big{)}+\alpha_{2}\big{(}H(t^{\prime}+1)-H(t^{\prime})\big{)}\Big{)}.$
  • the net delta over $\mathcal{G}$ to be the sum of the changes in $N(t)$ and $H(t)$ during the slots in $\mathcal{G}$, i.e.,
    $\sum_{t^{\prime}\in\mathcal{G}}\Big{(}\alpha_{1}\big{(}N(t^{\prime}+1)-N(t^{\prime})\big{)}+\alpha_{2}\big{(}H(t^{\prime}+1)-H(t^{\prime})\big{)}\Big{)}.$
Then, for proper choices of $\alpha_{1}$ and $\alpha_{2}$:
  • The net delta over $\mathcal{L}$ is at most $O(\ln^{2}\tau)+\alpha_{1}(\mathcal{A}_{\mathcal{L}}+\mathcal{J}_{\mathcal{L}})$ w.h.p. in $\tau$.
  • The net delta over $\mathcal{G}$ is at most $O(\ln^{2}\tau)+\alpha_{1}(\mathcal{A}_{\mathcal{G}}+\mathcal{J}_{\mathcal{G}})-\Omega(|\mathcal{G}|)$ w.h.p. in $\tau$.
Lemma 3.4 provides a symmetric high-probability bound on $N(t)+H(t)$ over the high-contention slots in an arbitrary interval. For the high-probability bound, in this case, we have that $N(t)+H(t)$ will decrease by $\Omega(|\mathcal{H}|)$, where $|\mathcal{H}|$ is the number of high-contention slots, up to the usual additional terms of jamming, packet injections, and a polylog term in terms of the interval length.
{restatable*}
[Net delta in contribution of $N(t)$ and $H(t)$ to potential over high contention slots]lemmaABhigh Let $\mathcal{I}$ be an arbitrary interval starting at time $t$ with length $|I|=\tau$. Let $\mathcal{H}$ be the set of slots in $\mathcal{I}$ during which $C(t)>C_{\rm high}$. Let $\mathcal{A}_{\mathcal{H}}$ be the number of packet arrivals in time slots in $\mathcal{H}$, and let $\mathcal{J}_{\mathcal{H}}$ be the number of jammed slots in $\mathcal{H}$,
Define the net delta over $\mathcal{H}$ to be the sum of the changes in $N$ and $H$ during the slots in $\mathcal{H}$, i.e.,
$\sum_{t^{\prime}\in\mathcal{H}}\Big{(}\alpha_{1}\big{(}N(t^{\prime}+1)-N(t^{\prime})\big{)}+\alpha_{2}\big{(}H(t^{\prime}+1)-H(t^{\prime})\big{)}\Big{)}.$
Then the net delta over $\mathcal{H}$ is at most $O(\ln^{3}\tau)+\alpha_{1}\mathcal{A}_{\mathcal{H}}-\Omega(|\mathcal{H}|)$ w.h.p. in $\tau$.
It is worth noting that the proofs of Lemmas 3.4 and 3.4 are technically involved. One of the reasons for this is that these lemmas contain our main applications of Theorems 4 and 5. This is necessary because the potential-function terms behave very differently in the three contention regimes—and because the adaptive adversary has the ability to change the contention in a slot on the fly.
Lemma 3.4 then collates Lemmas 3.4 and 3.4 to show that over an arbitrary interval of length $\tau$, it is either the case that almost all of the slots are low contention slots, or the first two terms decrease by $\Omega(\tau)$ with high probability (again, up to terms for packet insertions and jamming). Lemma 3.4 considers all slots, rather than only those of a particular contention regime. This lemma is the only one from this subsection that will be used later in the analysis, but the earlier lemmas in the subsection are necessary to build up to it.
{restatable*}
[Unless most slots have low contention, $\alpha_{1}N(t)+\alpha_{2}H(t)$ decreases]lemmaABnothigh Let $\mathcal{I}$ be an arbitrary interval of length $\tau>\Omega(1)$ with $\mathcal{A}$ packet arrivals and $\mathcal{J}$ jammed slots. With high probability in $\tau$, at least one of the following two conditions holds:
  • Less than $1/10$ of slots satisfy $C(t)\geq C_{\rm low}$.
  • $\alpha_{1}N(t)+\alpha_{2}H(t)$ decreases by $\Omega(\tau)-O(\mathcal{A}+\mathcal{J})$ over $\mathcal{I}$.
Additionally, $\alpha_{1}N(t)+\alpha_{2}H(t)$ increases by at most $O(\ln^{3}\tau+\mathcal{A}+\mathcal{J})$ w.h.p. in $\tau$.
Overview of Section 5.3.  Amortized behavior of $L(t)$.
This section analyzes $L(t)$’s behavior over intervals of length $|\mathcal{I}|=\tau=(1/c_{\tau})\cdot\max\{\frac{w_{\max}(t)}{\ln^{2}(w_{\max}(t))},N(t)^{1/2}\}$. The two main things that we want to show are that $L(t)$ does not increase by much, regardless of the contention regime, and that when there are many low-contention slots, $L(t)$ exhibits a substantial decrease.
Lemma 3.4 argues that a packet with large-enough window size is unlikely to have its window change by much during the interval. This lemma is instrumental when considering packets across an interval (notably in the proof of Lemma 3.4) as it means that their probability of listening also does not change by much.
Lemma 3.4 provides one of the main results of the section: a tail bound, and hence also a high probability bound, on how much $L$ increases over the interval regardless of contention. The proofs for both Lemmas 3.4 and 3.4 amount to arguing that an individual packet is not likely to listen to the channel too many times, which means that its window size also cannot change by very much. Because we are pessimistically counting the number of listens, the actual state of the channel does not appear in the proofs, and thus the number of jammed slots is irrelevant.
{restatable*}
[Bounds on the factor that a large window can grow/shrink]lemmaWchange Consider any packet during an interval $\mathcal{I}$ with $\tau=|\mathcal{I}|$. Let $Z$ satisfy $Z/\ln^{2}(Z)=\tau$. And let $W\geq w_{\rm min}$ be the initial size of the packet’s window. Let $W^{-}$ be the smallest window size the packet has while still active in the interval, and let $W^{+}$ be the biggest window size the packet achieves during the interval. Then for large enough choice of constants $w_{\rm min}$ and $c$ and any constant parameter $\gamma>0$ and $k\geq 2$:
If $W=\Theta(kZ)$, then
$\Pr\left[W^{+}\geq e^{\gamma}W\,\mbox{ or }\,W^{-}<W/e^{\gamma}\right]\leq 1/\tau^{\Theta(c\gamma\lg(\gamma k))}\,.$
{restatable*}
[Tail bound on increase in $L(t)$]lemmaLtail Consider an interval $\mathcal{I}$ with length $\tau=|\mathcal{I}|$ starting from time $t$ and ending at time $t^{\prime}=t+\tau$. Let $\mathcal{A}$ be the number arrivals during the interval. Then for for large-enough constant $c$ in the algorithm and any $k\geq 2$:
$\Pr\left[L(t^{\prime})\geq\Theta(\mathcal{A}+k\tau)\right]\leq 2^{-\Theta(c(\lg\tau\cdot\lg k+\lg^{2}k))}\,.$
Lemma 3.4 is the other main result of the section. This lemma says that as long as most slots have low contention, then $L$ decreases by $\Omega(\tau)$, minus the number of packet arrivals and jammed slots. The proof focuses on packets with large windows, i.e., window closes to $w_{\max}(t)$. The main idea of the proof is to give a high-probability lower bound on the number of times each such packet listens and hears silence as well as an upper bound on how many times the packet listens and hears noise. As long as the former is larger by a constant factor, the packet is likely to decreases its window size by a constant factor. Taking a union bound across packets is enough to conclude that all packets with large windows have their window sizes decrease, with high probability.
{restatable*}
[Mostly low contention implies decrease in $L(t)$]lemmaCterm Consider an interval $\mathcal{I}$ starting at $t$ of length $\tau$, where $\tau=(1/c_{\tau})\max\left\{L(t),\sqrt{N(t)}\right\}$. Let $t_{1}=t+\tau$, and let $\mathcal{A}$ and $\mathcal{J}$ denote the number of packet arrivals and jammed slots, respectively, over $\mathcal{I}$. Then, with high probability in $\tau$, either
  • $L(t_{1})\leq L(t)/d+O(\mathcal{A})$, where $d>1$ is a positive constant, or
  • At least a $1/10$-fraction of the slots $t^{\prime}$ in the interval $\mathcal{I}$ are either jammed or have contention $C(t^{\prime})\geq C_{\rm low}$.
Incorporating the fact that $\tau\geq L(t)/c_{\tau}$, it follows that if at least a $9/10$ fraction of slots in the interval have contention at most $C_{\rm low}$, then $(L(t_{1})-L(t))\leq O(\mathcal{A}+\mathcal{J})-\Omega(\tau)$.
Overview of Section 5.4.  Combining the analyses of $N(t)$, $H(t)$, and $L(t)$, to analyze $\Phi(t)$.
This section combines all three terms of the potential function to characterize the overall behavior of $\Phi(t)$. The key tools established in this section are Theorem 3.3 (stated previously in Section 3.3) and Theorem 3.4, which allow us to argue that the potential will decrease (most of the time) and that, when this fails to occur, the amount by which it increases is bounded. Specifically, consider a size-$\tau$ interval with $\mathcal{A}$ packet arrivals and $\mathcal{J}$ jammed slots. Theorem 3.3 shows that $\Phi(t)$ decreases by $\Omega(\tau)-O(\mathcal{A}+\mathcal{J})$ w.h.p. in $\tau$. Theorem 3.4 establishes tail bounds, proving that even when the high-probability bound of Theorem 3.3 fails, the probability that $\Phi(t)$ increases by more than $k\tau^{2}+O(\mathcal{A}+\mathcal{J})$ is less than $\frac{1}{\text{poly}(\tau)}\cdot{(1/2)}^{\Theta(\log^{2}k)}$.
Theorems 3.3 and 3.4 are the tools needed to fit our betting game, descussed next and in Section 5.5, and thereby argue that the potential is likely to decrease sufficiently across multiple intervals.
{restatable*}
[Tail bound on increase in $\Phi(t)$ over interval $\mathcal{I}$]theoremtailbound Consider an interval $\mathcal{I}$ of length $|\mathcal{I}|=\tau$ starting at time $t$ and ending at time $t^{\prime}$. Let $\mathcal{A}$ be the number of packet arrivals in $\mathcal{I}$. Then the probability that $\Phi$ increases by at least $\Theta(\mathcal{A})+\Theta(k\tau^{2})$ is at most $2^{-\Theta(c(\lg\tau\cdot\lg k+\lg^{2}k))}\leq(1/\tau^{\Theta(c)})\cdot 2^{-\Theta(\log^{2}k)}$, where $c$ is the constant parameter of the algorithm. That is,
$\Pr\left[(\Phi(t^{\prime})-\Phi(t))\geq\Theta(\mathcal{A})+\Theta(k\tau^{2})\right]\leq\left(\frac{1}{\tau^{\Theta(1)}}\right)\cdot 2^{-\Theta(\log^{2}k)}\ .$
Overview of Section 5.5. Using $\Phi(t)$ to prove throughput via a betting-game argument.
The analysis so far establishes progress guarantees over sufficiently large intervals in the form of Theorems 3.3 and 3.4. In this section, we show how to apply these theorems to give upper bounds on the potential over the execution with high probability in the total number of packets and the number of jammed slots.
Since the adversary is adaptive, we have to be careful in combining bounds across intervals. The adversary can use the results of earlier intervals in choosing new arrivals and jamming, which affects the size of later intervals. To reason about this process, we reframe it in a setting that resembles a random walk, which we describe below in a betting game. Our analysis of this game then allows us to analyze the implicit throughput (recall Section 1.1).
The Betting Game. We first summarize the betting game and then later relate it to the backoff process. The adversary corresponds to a bettor who makes a series of bets. Each bet has a size equal to the duration $\tau$. The bettor also has some amount of money, which is initially 00 dollars. When the bettor loses a bet, the bettor loses some money, and when the bettor wins, the bettor wins some money. (The amounts won or lost are specified below as a function of the size of the bet.) Additionally, at any time, the bettor may choose to receive a passive income. The passive income is added to the bettor’s wealth. The total amount of passive income taken, however, means that the bettor must play the game longer. The game begins when the bettor first takes some passive income, and the game does not end until either the bettor goes broke or the bettor has resolved bets totaling $\Omega(P)$ size, where $P$ is the passive income received, whichever comes first. The bettor’s goal is to complete the game without going broke. Importantly, although the bettor can always choose to take more passive income, doing so increases the total play time.
We set the details of the betting game to mirror the backoff process. Each bet corresponds to an interval. Passive income during a bet corresponds to the number of arrivals and jammed slots during the interval. Money corresponds to potential.
The bettor loses a size-$\tau$ bet with probability at least $1-\frac{1}{\text{poly}(\tau)}$. If the better loses the size-$\tau$ bet, it loses $\Theta(\tau)$ dollars. This loss corresponds to the high-probability event (in $\tau$) of Theorem 3.3. The bettor wins a size-$\tau$ bet with probability $O(1/\text{poly}(\tau))$. If the bettor wins the bet, it gets $\Theta(\tau^{2})$ dollars, plus $Y$ bonus dollars, where $Y$ is a random variable such that $\Pr[Y\geq k\tau^{2}]\leq\frac{1}{\text{poly}(\tau)}\cdot 2^{-\Theta(\log^{2}k)}$; these winnings correspond to tail bound of Theorem 3.4. (Of course, during each bet, the bettor can also gain passive income for arrivals and jammed slots.)
We pessimistically give the bettor the power to choose arbitrary bet sizes (subject to a minimum interval size, which itself is determined by $w_{\rm min}$), and the bettor is even allowed to place bets whose loss would cause the bettor to end with negative money. (In the actual backoff process, the interval sizes are dictated by the current state of the system, and not entirely under the control of the adversary.)
The rules of betting game are set pessimistically (in favor of the bettor) such that when the bettor wins, $\Phi(t)$ increases more slowly than the bettor’s wealth increases, and when the bettor loses, $\Phi(t)$ decreases at least as fast as the bettor’s wealth decreases. Therefore, this betting game stochastically dominates the potential function.
The takeaway is that at any point $t$, the bettor’s wealth is an upper bound on $\Phi(t)$. Because $\Phi(t)$ is an upper bound on the number of packets in the system, the bettor going broke corresponds to all packets succeeding. We thus obtain good implicit throughput, because there must either be many jammed slots or packet arrivals, or there must be many packets succeeding, leading to inactive slots.
Upper bounding the bettor’s maximum wealth/potential and showing $\Omega(1)$ implicit throughput.
In Lemma 3.4, we provide a high-probability upper bound on the bettor’s maximum wealth and the amount of time until it goes broke, which corresponds to there being no packets in the system.
{restatable*}
[The bettor loses the betting game] lemmaBGpassive Suppose the bettor receives $P$ dollars of passive income. Then with high probability in $P$, the bettor never has more than $O(P)$ dollars across the execution. Moreover, the bettor goes broke within $O(P)$ active slots, with high probability in $P$.
We briefly explain here how Lemma 3.4 implies implicit throughput. Consider a time horizon $t$, and suppose that the bettor has received $P=t/c$ dollars from passive income, for constant $c$ matching the big-$O$ of the lemma. Then from Lemma 3.4, with high probability in $t/c$, the bettor goes broke within $c\cdot t/c=t$ time; that is, there are no active packets at time $t$. We thus obtain the $\Omega(1)$ throughput result of Theorem 3.
Overview of Section 5.6.  Channel access/energy bounds.
In this section, we establish energy bounds. Two of the theorem statements, namely Theorems 3.4 and 3.4; the rest appear in Section 5.6. Theorems 3.43.4 are proved via properties of our potential function. (Several additional useful lemmas about the potential, not highlighted above, do appear in Section 5.5). So far, we have primarily motivated $\Phi(t)$ as a tool for proving throughput bounds, but $\Phi(t)$ also enables channel-access bounds.
Theorem 3.4 gives energy bounds in the finite case against an adaptive adversary. Specifically, if the stream has $N$ packets and $\mathcal{J}$ jammed slots, then w.h.p. each packet accesses the channel at most $\operatorname{polylog}(N+\mathcal{J})$ times. The proof structure is as follows: Our upper bound on $\Phi(t)$ immediately gives an upper bound on a packet’s maximum window size: $w_{\max}(t)=O(\text{poly}(\Phi(t))=O(\text{poly}(N+\mathcal{J}))$. Thus, if a packet accesses the channel too many times, then many of these accesses must have been listening during silent slots, so that the packet window can get smaller. However, by the structure of Low-Sensing Backoff, whenever a packet first chooses to listen, there is at least a $1/\operatorname{polylog}(N+\mathcal{J})$ probability that it also sends. Thus, after $\operatorname{polylog}(N+\mathcal{J})$ channel accesses when all other packets are silent, with high probability that packet has been transmitted.
{restatable*}
[Energy bound for finite case against adaptive adversary]theoremfiniteenergy Consider an input stream with $N$ packets and $\mathcal{J}$ jammed slots. Assume that the adversary is adaptive but not reactive. Then w.h.p. in $N+\mathcal{J}$, a packet accesses the channel at most $O(\log^{4}(N+\mathcal{J}))$ times.
The corresponding proof illustrates one subtle design choice of Low-Sensing Backoff, which leads to an easier energy analysis. Specifically, a given packet’s sending and listening probabilities are correlated: if a packet sends, then it has already decided to listen (but, of course, a packet can listen without deciding to send). We conclude by observing that, with an adaptive adversary, all packets have good channel-access bounds.
Theorem 15 gives an analogous result for an adversary that is both adaptive and reactive. By the very nature of a reactive adversary, there is no possibility of good per-packet bounds on channel accesses. (For example, a reactive adversary could target a specific packet and reactively jam whenever it sees this packet try to transmit.) However, interestingly, the amortized channel-access bounds are still good. This is because the reactive adversary only learns about sending on the channel and can react instantaneously; it does not learn whether a packet is listening in the current slot. Thus, a targeted packet can still reduce its window (as the other packets do) and it will succeed in sending unless the adversary does significant jamming. For example, consider the special case where the targeted packet is the only packet remaining. Then, unless the adversary (which does not sense when a packet will listen) jams a large number of slots, this packet will correctly back on and transmit successfully.
Theorems 16 and 17 generalize Theorems 3.4 and 15 to the adversarial-queuing setting with granularity $S$ and sufficiently small arrival rate $\lambda$. The main tool is Lemma 13 , which allows us to transform the adversarial queuing case into finite instances that are not very large. Theorem 3.4 applies to infinite streams with arbitrary arrivals.
{restatable*}
[Channel access bounds for infinite case against adaptive and reactive adversaries]theoreminfiniteenergy Suppose that up until time $t$ there have been $N_{t}$ packet arrivals and $\mathcal{J}_{t}$ jammed slots.
  • Consider an adaptive adversary that is not reactive. Then w.h.p. in $\mathcal{J}_{t}+N_{t}$, each packet makes $O(\log^{4}(\mathcal{J}_{t}+N_{t}))$ channel accesses before time $t$.
  • Consider and adaptive adversary that is reactive. Then w.h.p. in $\mathcal{J}_{t}+N_{t}$, a particular packet accesses the channel at most $O((\mathcal{J}_{t}+1)\log^{3}(N_{t}+\mathcal{J}_{t})+\log^{4}(N_{t}+\mathcal{J}_{t}))$ times. Moreover, the average number of channel accesses is $O((\mathcal{J}_{t}/N_{t}+1)\log^{4}(N_{t}+\mathcal{J}_{t}))$.
4. Conclusion
We have provided a simple contention-resolution algorithm that achieves constant throughput with full energy efficiency (i.e., low sending and listening complexity), despite a powerful jamming adversary. This resolves in the affirmative two open questions about whether full-energy efficiency is possible at all in the popular ternary-feedback model, and whether it remains possible in the presence of jamming.
Acknowledgements. This work is supported by NSF awards CCF-1918989, CCF-2106759, CNS 2210300, and CCF 2144410.
Appendix A
In this appendix, we start by providing the full proofs for our analysis in Section 5. The only exceptions are the proofs in Section 5.1 (Preliminaries) and for the two observations (Observations 7 and 8) in Section 5.2, which we defer until Appendix B, in order to streamline the exposition. Additional related results from the literature are summarized Section 6. Finally, we wrap up with a conclusion and a discussion of future work in Section 4.
5. Analysis
Our analysis is presented in the same order as discussed in our overview of the proof organization in Section 3.4.
5.1. Preliminaries
Fact 1 ().
The following inequalities hold.
  1. (a)
    For any $10\leq x<$ 1, $1-x\geq e^{-x/(1-x)}$.
  2. (b)
    For any $10\leq x<$ 1, $\ln(1+x)=x-\sum_{j=2}^{\infty}(-x)^{j}/j$.
  3. (c)
    For any $20\leq x<1/$ 2, $1+2x\geq e^{x}$.
  4. (d)
    For any $x$, $1-x\leq e^{-x}$.
For any fixed slot $t$, let $p_{\rm suc}(t)$ denote the probability that some message is successful in slot $t$. We state the following result relating $C(t)$ and $p_{\rm suc}(t)$; this has appeared in similar forms in (richa:jamming4, ; BenderFiGi19, ; DBLP:conf/spaa/AgrawalBFGY20, ; BenderFiGi16, ; ChangJiPe19, ).
Lemma 1 (Probability of success as a function of contention).
For any unjammed slot $t$, where $w_{u}(t)\geq 2$ for all packets, the following holds:
$\frac{C(t)}{e^{2C(t)}}\leq p_{\rm suc}(t)\leq\frac{2C(t)}{e^{C(t)}}.$
Note that any packet’s window sizes is always at least $2$ under Low-Sensing Backoff, so this lemma is always applicable for our algorithm.
Let $p_{\rm emp}(t)$ denote the probability that slot $t$ is empty, and let $p_{\rm noi}(t)$ denote the probability that slot $t$ is noisy. We have the following lower bounds on these probabilities (the proofs are deferred to the Appendix).
Lemma 2 (Probability of an empty slot as a function of contention).
For any unjammed slot $t$, where $w_{u}(t)\geq 2$ for all packets, the following holds:
$e^{-2C(t)}\leq p_{\rm emp}(t)\leq e^{-C(t)}.$
Lemma 3 (Probability of a noisy slot as a function of contention).
For any unjammed slot $t$, where $w_{u}(t)\geq 2$ for all packets, $p_{\rm noi}(t)\geq 1-\frac{2C(t)}{e^{C(t)}}-\frac{1}{e^{C(t)}}$.
Let $p_{{\rm suc}|{\rm good}}$ be a lower bound on the probability of success when contention is good in a slot. Let $p_{{\rm emp}|{\rm low}}$ be a lower bound on the probability of an empty slot when contention is low in a given slot. Let $p_{{\rm noi}|{\rm high}}$ be a lower bound on the probability of a noisy slot when contention is high in slot $t$. Finally, let $p_{{\rm emp}|{\rm good}}$ be a lower bound of the probability that a slot is silent given that $C(t)\leq C_{\rm high}$. A constant lower bound can be established for each event, so we set all of these quantities to $\Theta(1)$, which follows directly from Lemmas 1, 2, and 3.
Throughout this paper, we often want to use Chernoff bounds to bound some random process over various intervals. The problem is that the probability distributions vary in each time step, i.e., there are different probabilities of listening, sending, etc., depending on the contention in each slot as well as the window sizes of the packets that make up that contention. On top of that, we want to apply different analyses depending on whether contention is good, low, or high, and the number of slots within each contention category is something that is also determined by the adversary adaptively. To simplify enormously (and in some cases, enable) these Chernoff-style arguments, we use the following theorems from (KuszmaulQi21, ), which generalize Chernoff bounds using a martingale style analysis, allowing the probability distribution in one time step to depend on the outcomes of previous time steps.
Theorem 4 (Corollary 11 from (KuszmaulQi21, )).
Suppose that Alice constructs a sequence of random variables $X_{1},\ldots,X_{n}$, with $X_{i}\in[0,c]$, $c>0$, using the following iterative process: once the outcomes of $X_{1},\ldots,X_{i-1}$ are determined, Alice then selects the probability distribution $\mathcal{D}_{i}$ from which $X_{i}$ will be drawn; $X_{i}$ is then drawn from distribution $\mathcal{D}_{i}$. Alice is an adaptive adversary in that she can adapt $\mathcal{D}_{i}$ to the outcomes of $X_{1},\ldots X_{i-1}$. The only constraint on Alice is that $\sum_{i}\mathbb{E}[X_{i}|D_{i}]\leq\mu$, that is, the sum of the means of the probability distributions $\mathcal{D}_{1},\ldots,\mathcal{D}_{n}$ must be at most $\mu$.
If $X=\sum_{i}X_{i}$, then for any $\delta>0$,
$\Pr[X\geq(1+\delta)\mu]\leq\exp{\left(-\frac{\delta^{2}\mu}{(2+\delta)c}\right)}.$
If the value of $\mu$ is not fixed in advance (but is a bounded and well-defined result of Alice’s choices, and at least $1$), then the same result is achieved with a slightly weaker bound:
$\Pr[X\geq 2(1+\delta)\mu]\leq O\left(\exp{\left(-\frac{\delta^{2}\mu}{(2+\delta)c}\right)}\right).$
Proof.
When $\mu$ is known in advance, this theorem is proved as Corollary 11 in (KuszmaulQi21, ). When $\mu$ is not known in advance, we can construct a sequence of such games for $\mu=1,2,4,\ldots$. Each of these games uses the same sequence of distributions from Alice. If Alice’s choices satisfy the $\mu$ for that specific game, then the above tail bound holds. Since the error probability decreases exponentially with $\mu$, we can take a union bound over all the games for which Alice’s choices satisfy the requirement, with the error dominated by the smallest $\mu$. ∎
Theorem 5 (Corollary 16 from (KuszmaulQi21, )).
Suppose that Alice constructs a sequence of random variables $X_{1},\ldots,X_{n}$, with $X_{i}\in[0,c],c>0$, using the following iterative process. Once the outcomes of $X_{1},\ldots,X_{i-1}$ are determined, Alice then selects the probability distribution $\mathcal{D}_{i}$ from which $X_{i}$ will be drawn; $X_{i}$ is then drawn from distribution $\mathcal{D}_{i}$. Alice is an adaptive adversary in that she can adapt $\mathcal{D}_{i}$ to the outcomes of $X_{1},\ldots X_{i-1}$. The only constraint on Alice is that $\sum_{i}\mathbb{E}[X_{i}|D_{i}]\geq\mu$, that is, the sum of the means of the probability distributions $\mathcal{D}_{1},\ldots,\mathcal{D}_{n}$ must be at least $\mu$.
If $X=\sum_{i}X_{i}$, then for any $\delta>0$,
$\Pr[X\leq(1-\delta)\mu]\leq\exp{\left(-\frac{\delta^{2}\mu}{2c}\right)}.$
If the value of $\mu$ is not fixed in advance (but is a bounded and well-defined result of Alice’s choices, and at least $1$), then the same result is achieved with a slightly weaker bound:
$\Pr[X\leq(1-\delta)\mu/2]\leq O\left(\exp{\left(-\frac{\delta^{2}\mu}{2c}\right)}\right).$
Proof.
When $\mu$ is known in advance, this theorem is proved as Corollary 16 in (KuszmaulQi21, ). When $\mu$ is not known in advance, we can construct a sequence of such games for $\mu=1,2,4,\ldots$. Each of these games uses the same sequence of distributions from Alice. If Alice’s choices, satisfy the $\mu$ for that specific game, then the above tail bound holds. Since the error probability decreases exponentially with $\mu$, we can take a union bound over all the games for which Alice’s choices satisfy the requirement, with the error dominated by the smallest $\mu$. ∎
5.2. $N(t)+H(t)$: over single slots and intervals, when contention is low, high, and good
Analyzing $N(t)$: over single slots and intervals, when $C(t)$ is good.
We ease into our analysis by starting with the first term of our potential function, $N(t)$, in the special case where contention is constant. Lemma 6 analyses the expected decrease in $N(t)$ on a per-slot basis when contention is good.
Lemma 6 ().
Consider a slot where contention is good, i.e., $C(t)\in[C_{\rm low},C_{\rm high}]$, and the number of packet arrivals is $\mathcal{A}$. Then, $N(t)$ decreases by at least $p_{{\rm suc}|{\rm good}}-\mathcal{A}=\Theta(1)-\mathcal{A}$ in expectation.
Proof.
By Lemma 1, $p_{{\rm suc}|{\rm good}}(t)\geq C(t)\,e^{-2C(t)}\geq C_{\rm low}\,e^{-2C_{\rm high}}=\Theta(1)$, which is thus the expected amount by which $N(t)$ decreases. Packet arrivals increase $N(t)$ by $1$ per packet arrival. ∎
How $H(t)$ changes in expectation on a per-slot basis.
Observation 7 shows that, in a noisy slot, $H(t)$ decreases in expectation by $\Omega(C(t))$. Conversely, Observation 8 shows that, in a silent slot, $H(t)$ increases in expectation by $O(C(t))$. Given that these next two observations only offer intuition, their accompanying analysis is deferred to Section Proofs for Section 5.2 of Appendix B.
Observation 7 ().
Assume no packet arrivals and no jamming. Then, if slot $t$ is noisy, we have $\mathbb{E}[H(t+1)~{}|~{}$t $\mbox{~{}is noisy}]\leq H(t)-\frac{C(t)}{2c}$.
Observation 8 ().
Assume no packet arrivals and no jamming. Then, if slot $t$ is silent, we have $\mathbb{E}[H(t+1)~{}|~{}$t $\mbox{~{}is silent}]\leq H(t)+\frac{2C(t)}{c}$.
How one packet listening in slot $t$ changes $H(t)$.
Lemma 3.4 illustrates how $H(t)$ increases or decreases due to a specific packet listening to an empty or noisy slot.
[…]
Proof.
Packet $u$’s contribution to $H(t)$ is $1/\ln(w_{u})$. After listening to a silent slot, $u$’s window size becomes $w_{u}/(1+\frac{1}{c\ln w_{u}})$. Thus the change to its term in $H(t)$ is
$\frac{1}{\ln w-\ln(1+\frac{1}{c\ln w})}-\frac{1}{\ln w}$ $={\textstyle\Theta\big{(}\ln\big{(}1+\frac{1}{c\ln w}\big{)}/\ln^{2}w\,\big{)}}$
$=\Theta(1/(c\ln w)/\ln^{2}w)=\Theta(1/(c\ln^{3}w))\,.$
Similarly, if the slot is noisy, $u$’s window size becomes $w_{u}\cdot(1+1/c\ln w_{u})$. Thus, the change to $u$’s contribution to $B$ is
$\frac{1}{\ln w+\ln(1+1/(c\ln w))}-\frac{1}{\ln w}$ $=-\Theta(\ln(1+1/(c\ln w)/\ln^{2}w))=-\Theta(1/(c\ln^{3}w))\,.$
Net behavior of $N(t)+H(t)$ during low and good contention slots.
We now bound the behavior of $N(t)$ and $H(t)$ during non high contention slots. Specifically, we show that these terms do not increase too much in low-contention slots and that they decrease in good-contention slots, with high probability.
[…]
Proof.
Since we are only considering slots in the low contention regime and good contention regime, $H(t)$ will generally not be decreasing significantly. Instead, our line of argument will be to show that $N(t)$ will decrease over slots with good contention, overwhelming any increase from $H(t)$. And in slots with low contention, $N(t)$ will decrease sufficiently to counterbalance any increase from $H(t)$ (up to an additive $O(\ln^{2}\tau)$ term).
In this lemma, we are going to analyze the sending attempts and the listening attempts separately. When packet transmissions are successful, they reduce $N(t)$; when listen events are successful, they (may) increase $H(t)$ (if the slot is empty). We will show that in each case, the change in $N(t)$ or $H(t)$ is proportional to the sum of contention in the slots being considered, with high probability. By choosing $\alpha_{1}$ sufficiently large with respect to $\alpha_{2}$, we can ensure that reduction of $N(t)$ due to successful transmissions dominates the increase of $H(t)$ due to listening.
The adversary, being adaptive, has some control over which slots packets are listening/sending in, and which slots have non-high contention. Throughout this proof, we think of the process as a game where in each slot the adversary adaptively determines: (i) what the contention is, and (ii) what the window sizes are that lead to this contention. Thus, we allow the adversary some additional power.
Upper bound for listening: bounding the increase to $H$.
First, we focus on listening attempts. In this proof, we are going to assume (the worst case) that every listening attempt hears silence and thus decreases its window size, increasing the $H(t)$ term. As we have seen in Lemma 3.4, if a packet with window size $w$ listens to a silent slot, then its contribution to $H(t)$ increases by $\Theta(1/\ln^{3}w)$ additively.
Let $X_{1},\ldots X_{n}$ be random variables associated with listen attempts in $\mathcal{I}$ during slots in either $\mathcal{L}$ or $\mathcal{G}$. (The same analysis holds for both of these subsets of slots.) For random variable $X_{j}$, if the associated listener has window size $w$ and the listen succeeds (i.e., the packet does listen), we define $X_{j}=1/\ln^{3}w$; otherwise, $X_{j}=0$. Notice that each $X_{j}$ is in the range $[0,1]$, and $X=\sum X_{i}$ upper bounds the total change, within constant factors, of $H(t)$ during the slots being analyzed.
Let $\mu_{x}={\rm E}\left[X\right]$. While the adversary can adaptively determine the contention in each slot, we can still express $\mu_{x}$ as a function of the selected contention. Let $X^{t}$ be the sum of the $X_{j}$ associated with time slot $t$. Since ${\rm E}\left[X_{j}\right]=\Theta(1/w)$, if the corresponding packet has window size $w$, we know that ${\rm E}\left[X^{t}\right]=\Theta(C(t))$. Thus, if we are analyzing the slots in $\mathcal{L}$: $\mu_{x}=\Theta(\sum_{t\in\mathcal{L}}C(t))$ (i.e., the sum of the contention of low contention slots). If we are analyzing the slots in $\mathcal{G}$: $\mu_{x}=\Theta(\sum_{t\in\mathcal{G}}C(t))$ (i.e., the sum of the contention of good contention slots).
We can now apply Theorem 4 to show that $X$ is close to $\mu_{x}$. Notice that the precise distribution for each $X_{i}$ is determined adaptively by the adversary, e.g., by affecting which slots are high or low contention. However, each listen attempt is independent.
Theorem 4 shows that:
$\Pr[X\geq(1+\delta)\mu_{x}]\leq\exp{\left(-\frac{\delta^{2}\mu}{(2+\delta)}\right)}.$
If $\mu_{x}$ is $O(\log{\tau})$, we set $\delta=\Theta(\ln{\tau})$, and conclude that with high probability in $\tau$, $X\leq O(\ln^{2}(\tau))$. Otherwise, if $\mu_{x}$ is $\Omega(\ln(\tau))$, we can choose $\delta=O(1)$, and observe that with high probability in $\tau$, $X\leq O(\mu_{x})$.
We conclude, then, that the following two claims hold:
  • The change in $H(t)$ during $\mathcal{L}$ is at most $O(\sum_{t\in\mathcal{L}}C(t)+\ln^{2}(\tau))$, with high probability in $\tau$.
  • The change in $H(t)$ during $\mathcal{G}$ is at most $O(\sum_{t\in\mathcal{G}}C(t)+\ln^{2}(\tau))$, with high probability in $\tau$.
Lower bound for sending: lower bounding the decrease to $N$.
Next, we similarly consider sending attempts. We will temporarily ignore jamming, and determine the number of transmissions that would have been successful if there were no jamming. Let $Y_{1},\ldots Y_{n}$ be random variables associated with time slots in $\mathcal{L}$ or in $\mathcal{G}$. (As before, we analyze these two regimes simultaneously.) Define each $Y_{j}=1$ if there is a successful transmission in the slot associated with $Y_{j}$, and $Y_{j}=0$ otherwise. Notice that each $Y_{j}$ is in the range $[0,1]$, and $Y=\sum Y_{i}$. Let $\mu_{y}={\rm E}\left[Y\right]$.
Again, we relate $\mu_{y}$ to the contention values in the relevant slots. For a given slot $t$, the probability $\Pr(Y_{j}=1)=\Theta(C(t)/e^{2C(t)})=\Theta(C(t))$—this follows from Lemma 1, because we are considering only slots where $C(t)\leq C_{\rm high}=O(1)$. Thus, when analyzing $\mathcal{L}$: $\mu_{y}=\Theta(\sum_{t\in\mathcal{L}}C(t))$ (i.e., the sum of the contention of the low contention slots). When analyzing $\mathcal{G}$: $\mu_{y}=\Theta(\sum_{t\in\mathcal{G}}C(t))$ (i.e., the sum of the contention of the good contention slots).
We can now apply Theorem 5 to show that $Y$ is close to its expectation. As with $X_{i}$ above, the precise distribution for each $Y_{i}$ is determined adaptively by the adversary, e.g., by affecting which slots are high or low contention. However, the $Y_{i}$ are independent.
Theorem 5 shows that:
$\Pr[X\leq(1-\delta)\mu]\leq\exp{\left(-\frac{\delta^{2}\mu}{2}\right)}.$
If $\mu_{y}$ is $O(\ln{\tau})$, then we trivially conclude that $Y\geq\mu_{y}-O(\ln(\tau))$. Otherwise, if $\mu_{y}$ is $\Omega(\ln(\tau))$, we can choose $\delta=\Theta(1)$, and observe that with high probability in $\tau$, $Y\geq\Theta(\mu_{y})$.
So far, we have analyzed the number of successful packet transmissions, ignoring jamming. In fact, up to $\mathcal{J}_{\mathcal{L}}$ or $\mathcal{J}_{\mathcal{G}}$ slots may be jammed, and each jammed message reduces the number of successful transmissions by at most one. We conclude, then, that the following two claims hold:
  • The change in $N(t)$ during $\mathcal{L}$ is at most $O(\ln(\tau))-\Omega(\sum_{t\in\mathcal{L}}C(t))+\mathcal{A}_{\mathcal{L}}+\mathcal{J}_{\mathcal{L}}$, with high probability in $\tau$.
  • The change in $N(t)$ during $\mathcal{G}$ is at most $O(\ln(\tau))-\Omega(\sum_{t\in\mathcal{G}}C(t))+\mathcal{A}_{\mathcal{G}}+\mathcal{J}_{\mathcal{G}}$, with high probability in $\tau$.
Taking a union bound over the above four bulleted claims (for each of the two terms, there is a claim for each of the two contention regimes), all of them hold with high probability in $\tau$.
Combining sending and listening.
Combining our previous claims for sending and listen, the net delta over $\mathcal{L}$ is at most:
$\alpha_{2}O\big{(}\sum_{t\in\mathcal{L}}C(t)\big{)}-\alpha_{1}\Omega\big{(}\sum_{t\in\mathcal{L}}C(t)\big{)}+(\alpha_{1}+\alpha_{2})O(\ln^{2}(\tau))+\alpha_{1}(\mathcal{A}_{\mathcal{L}}+\mathcal{J}_{\mathcal{L}}),$
with high probability in $\tau$. Analogously, for $\mathcal{G}$, we have shown that the net delta over $\mathcal{G}$ is:
$\alpha_{2}O\big{(}\sum_{t\in\mathcal{G}}C(t)\big{)}-\alpha_{1}\Omega\big{(}\sum_{t\in\mathcal{G}}C(t)\big{)}+(\alpha_{1}+\alpha_{2})O(\ln^{2}(\tau))+\alpha_{1}(\mathcal{A}_{\mathcal{G}}+\mathcal{J}_{\mathcal{G}}),$
with high probability in $\tau$.
Hence, by choosing $\alpha_{1}$ to be a constant sufficiently larger multiplicatively than $\alpha_{2}$, the decrease in $N$ due to transmissions overwhelms the (possible) increase in $H$ due to listening. We conclude that (with high probability in $\tau$), the net delta over $\mathcal{L}$ is at most
$(\alpha_{1}+\alpha_{2})O(\ln^{2}(\tau))-O\big{(}\alpha_{1}\sum_{t\in\mathcal{L}}C(t))\big{)}+\alpha_{1}(\mathcal{A}_{\mathcal{L}}+\mathcal{J}_{\mathcal{L}}),$
implying that the net delta over $\mathcal{L}$ is at most $O(\ln^{2}\tau)+\alpha_{1}(\mathcal{A}_{\mathcal{L}}+\mathcal{J}_{\mathcal{L}})$ w.h.p. in $\tau$.
For $\mathcal{G}$, we observe something stronger: since $C(t)\geq C_{\rm high}\geq\Omega(1)$ in each slot, we know that $\sum_{t\in\mathcal{G}}C(t)=\Omega(|\mathcal{G}|)$. Thus, for $\mathcal{G}$, we conclude that (with high probability in $\tau$), the net delta over $\mathcal{G}$ is at most
$(\alpha_{1}+\alpha_{2})O(\ln^{2}(\tau))-\alpha_{1}\Omega(|\mathcal{G}|)+\alpha_{1}(\mathcal{A}_{\mathcal{L}}+\mathcal{J}_{\mathcal{L}}),$
implying that the net delta over $\mathcal{G}$ is at most $O(\ln^{2}\tau)+\alpha_{1}(\mathcal{A}_{\mathcal{G}}+\mathcal{J}_{\mathcal{G}})-\Omega(|\mathcal{G}|)$ w.h.p. in $\tau$. ∎
Cumulative behavior of $N(t)+H(t)$ during high-contention slots.
The previous lemma considered the slots of the interval where the contention was not high; the next lemma considers the slots of the interval where the contention is high. In this case, the net delta decreases due to decreases in $H(t)$.
[…]
Proof.
For our line of argument, we do not require $N(t)$ to decrease (and indeed packets do not often succeed when contention is high), and so simply bound the increase to $N(t)$ by $\mathcal{A}_{\mathcal{H}}$ due to packet arrivals. We focus for the rest of the proof on the change in $H(t)$ over $\mathcal{H}$.
When the contention is high, a large constant fraction of slots (in expectation) in $\mathcal{H}$ are noisy, and in each noisy slot, $H(t)$ decreases. Some of the slots (up to a constant fraction in expectation) will be empty, however, and any packets that listen in such slots will decrease their window size, increasing $H(t)$ (rather than decreasing it). We will show that the decrease is sufficiently more than the increase to achieve the desired result.
Since the adversary is adaptive, it has some control over which slots are in this contention class, which slots have more listeners than others, and which slots are more likely to be empty than others. Throughout this proof, we think of this as a game where in each slot the adversary adaptively determines: (i) what the contention is, and (ii) what the window sizes are that lead to this contention. (Thus we allow the adversary some additional power.) We analyze the change in $H(t)$ over $O(\ln(\tau))$ contention classes: $C_{\rm high},{C_{\rm high}+1},\ldots,\Theta(\ln(\tau))$. Fix $C$ to be one of those contention classes, and define $\mathcal{H^{\prime}}$ to be the slots in $\mathcal{H}$ with contention in the range $(C,C+1)$. (The last class includes all slots with contention $\Theta(\ln(\tau))$ or larger.)
We separately bound the number of listens in noisy slots (which decrease $H(t)$) and in empty slots (which increase $H(t))$. Let $\mathcal{E}$ be the subset of slots in $\mathcal{H^{\prime}}$ that are empty; let $\mathcal{F}$ be the subset of slots in $\mathcal{H^{\prime}}$ that are noisy.
By Lemma 3.4, if a packet with window size $w$ listens to a silent slot, then its contribution to $H(t)$ increases by $\Theta(1/\ln^{3}w)$ additively; if it listens to a noisy slot, then its contribution to $H(t)$ decreases by $\Theta(1/\ln^{3}w)$. Notice that jamming actually increases the likelihood that the slot is noisy (to a certainty), and so only helps to decrease $H(t)$.
Analyzing noisy slots.
Let $X_{1},\ldots X_{n}$ be random variables associated with listen attempts in $\mathcal{H}$ during slots in $\mathcal{F}$. For random variable $X_{j}$, if the associated listener has window size $w$ and the listen succeeds, we define $X_{j}=1/\ln^{3}(w)$; otherwise, $X_{j}=0$. Since the slots are (by definition) noisy, $H(t)$ will decrease by $\Omega(X_{j})$ by Lemma 3.4.
There remains one challenge with analyzing the $X_{j}$: we have restricted our attention to noisy slots, and in noisy slots it is more likely that a given packet has listened (and then chosen to send). Luckily, knowing that a slot is noisy only increases the likelihood of listening, so we can conclude that $\Pr(X_{j}\neq 0)\geq\Theta(\ln^{3}(w_{u})/w_{u})$. [^8] Technically, we can see that the probability of a noisy slot, given that a packet is listening, is at least the probability of an (unconditioned) noisy slot, and so Bayes law says that the probability of listening in a noisy slot is a least the probability of (unconditioned) listening. Thus we define a new set of independent $\hat{X_{j}}$ random variables where $\hat{X_{j}}=1/\ln^{3}(w)$ with probability $\Theta(\ln^{3}(w_{u})/w_{u})$ which are stochastically dominated by the original $X_{j}$ random variables.
Notice that $\hat{X}=\sum\hat{X_{j}}$ determines the total change, within constant factors, of $H(t)$ during the slots being analyzed. Let $\mu_{x}={\rm E}\left[\hat{X}\right]$. Since ${\rm E}\left[\hat{X_{j}}\right]=\Theta(1/w)$, if the packet has window size $w$, we know that $\mu_{x}=\Theta(\sum_{t\in\mathcal{F}}C(t))$ (i.e., the sum of the contention of slots in $\mathcal{F}$).
Since each $\hat{X_{j}}$ is in the range $[0,1]$ and they are independent, we can now apply Theorem 5 to show that $\hat{X}$ is close to $\mu_{x}$. Notice that the precise distribution for each $\hat{X_{j}}$ is determined adaptively by the adversary, e.g., by affecting which slots are high or low contention. However, the specific coin flips for each listen attempt are independent.
Theorem 5 shows that:
$\Pr[\hat{X}\leq(1-\delta)\mu_{x}]\leq\exp{\left(-\frac{\delta^{2}\mu_{x}}{2}\right)}.$
If $\mu_{x}$ is $O(\log{\tau})$, then we trivially conclude that $\hat{X}\geq\mu_{x}-O(\ln(\tau))$. Otherwise, if $\mu_{x}\geq\Omega(\ln(\tau))$, we can choose $\delta=\Theta(1)$, and observe that with high probability in $\tau$, $\hat{X}$ is $\Theta(\mu_{x})$.
Since the $\hat{X_{j}}$ stochastically dominate the $X_{j}$, we conclude, then, that $H(t)$ during $\mathcal{F}$ decreases by at least $\Omega(\sum_{t\in\mathcal{F}}C-\ln^{2}(\tau))$, with high probability in $\tau$.
Analyzing empty slots.
We can now apply the same analysis to the empty slots. Let $X_{1},\ldots X_{n}$ be random variables associated with listen attempts in $\mathcal{H}$ during slots in $\mathcal{E}$. For random variable $X_{j}$, if the associated listener has window size $w$ and the listen succeeds, we define $X_{j}=1/\ln^{3}w$; otherwise, $X_{j}=0$. Since the slots are (by definition) empty, $H(t)$ will increase by $\Omega(X_{j})$.
Knowing that a slot is empty only decreases the likelihood of listening, so we can conclude that $\Pr(X_{j}\neq 0)\leq\Theta(\ln^{3}(w_{u})/w_{u})$. [^9] Technically, we can see that the probability of an empty slot, given that a packet is listening, is less than the probability of an (unconditional) empty slot, and so Bayes law says that the probability of listening in an empty slot is less than the probability of (unconditional) listening. Thus we define a new set of independent $\hat{X_{j}}$ random variables where $X_{j}=1/\ln^{3}w$ with probability $\Theta(\ln^{3}(w_{u})/w_{u})$ which stochastically dominate (from above) the original $X_{j}$ random variables.
Notice that each $\hat{X_{j}}$ is in the range $[0,1]$, and $\hat{X}=\sum\hat{X_{i}}$ bounds the total change, within constant factors, of $H(t)$ during the slots being analyzed. Let $\mu_{x}={\rm E}\left[\hat{X}\right]$. Since ${\rm E}\left[\hat{X_{j}}\right]=\Theta(1/w)$, if the packet has window size $w$, we know that $\mu_{x}=\Theta(\sum_{t\in\mathcal{E}}C(t))$ (i.e., the sum of the contention of slots in $\mathcal{E}$).
We can now apply Theorem 4 to show that $\hat{X}$ is close to $\mu_{x}$. Theorem 4 shows that:
$\Pr[\hat{X}\geq(1+\delta)\mu_{x}]\leq\exp{\left(-\frac{\delta^{2}\mu_{x}}{(2+\delta)}\right)}.$
If $\mu_{x}<O(\log{\tau})$, we set $\delta=\Theta(\ln{\tau})$, and conclude that with high probability in $\tau$, $\hat{X}\leq O(\ln^{2}(\tau))$. Otherwise, if $\mu_{x}\geq\Omega(\ln(\tau))$, we can choose $\delta=O(1)$, and observe that with high probability in $\tau$, $\hat{X}\leq O(\mu_{x})$.
Since the $\hat{X_{j}}$ stochastically dominate the $X_{j}$, we conclude, then, that $H(t)$ during $\mathcal{E}$ increases by at most $O(\sum_{t\in\mathcal{E}}(C+1)+\ln^{2}(\tau))$, with high probability in $\tau$, assuming that we are not considering the “last” contention class containing all the slots with contention $\Theta(\ln\tau)$ or higher (which we will handle separately).
Relating the noisy and empty slots.
We now need to combine our analysis of noisy and empty slots, and then sum across all the different contention classes. The remaining observation is that a large constant fraction of the slots are noisy. The probability that a slot $t$ is noisy is at least $(1-e^{-\Theta(C(t)})\geq 3/4$ (for appropriate choice of $C_{\rm high}$), whereas the probability that a lost $t$ is empty is at most $1/4$. So most of the slots in each class will be noisy.
There are now three cases to deal with: (i) $C<O(\ln(\tau))$ and the number of slots is small, i.e., $\mathcal{H^{\prime}}<O(\ln(\tau))$; (ii) $C<O(\ln(\tau))$ and the number of slots is not small, i.e., $\mathcal{H^{\prime}}>\Omega(\ln(\tau))$; $C>\Omega(\ln(\tau))$.
In the first case, if $\mathcal{H^{\prime}}<O(\ln(\tau))$, we will simply assume the worst case, i.e., all the slots are empty. As we have shown above, this implies that with high probability in $\tau$, the term $H(t)$ increases by at most $O(C\ln(\tau)+\ln^{2}(\tau))=O(\ln^{2}(\tau))$ during these slots. Since there are only $O(\ln(\tau))$ different contention classes, this will increase $H(t)$ by at most $O(\ln^{3}(\tau))$ throughout the entirety of $\mathcal{H}$.
In the second case, if $\mathcal{H^{\prime}}>\Omega(\ln(\tau))$: if there are $k$ slots in $\mathcal{H^{\prime}}$, then we expect $(3/4)k$ of them to be noisy; we can again use Theorem 5 to show that at least $(2/3)k$ of them are noisy, and at most $(1/3)k$ of them are empty (e.g., choosing $\delta=1/9$), with high probability in $\tau$. Thus, with high probability in $\tau$, $H(t)$ will decrease by at least $\Omega(k((2/3)C-(1/3)(C+1))-\ln^{2}(\tau))=\Omega(|\mathcal{H^{\prime}}|-\ln^{2}(\tau))$. (Notice that jammed slots only increase the number of slots that are noisy in the above analysis.)
In the third case, for the “last” contention class where $C>\Omega(\ln\tau)$, we observe that a slot is empty with probability at most $e^{-\Theta(C(t)}\leq 1/\tau^{\Theta(1)}$. Thus, with high probability (by a union bound) every slot in $\mathcal{H^{\prime}}$ is noisy. Thus, by the analysis above, we know that with high probability in $\tau$, $H(t)$ decreases by at least $\Omega(|\mathcal{H^{\prime}}\ln(\tau)|-\ln^{2}(\tau))$.
Overall, there are at least $|\mathcal{H}|-O(\ln^{2}(\tau))$ slots in the second and third cases. So summing up over all the cases (and taking a union bound over everything), we conclude that with high probability in $\tau$, $H(t)$ decreases by at least $\Omega(|\mathcal{H}|-\ln^{3}(\tau))$. ∎
Combined analysis of $N(t)+H(t)$ over an interval for all slots.
The next lemma aggregates Lemmas 3.4 and 3.4 to characterize the behavior of $N(t)+H(t)$ over all slots in an arbitrary interval. Specifically, Lemma 3.4 shows that, for an arbitrary interval of length $\tau$, it is either the case that almost all of the slots are low contention slots, or $N(t)+H(t)$ decreases by $\Omega(\tau)$ with high probability, blunted by the number of packet injections and jammed slots. This lemma addresses all slots, rather than being limited to those of a particular contention regime.
[…]
Proof.
Let $\mathcal{L}$ be the slots in $\mathcal{I}$ with contention $<C_{\rm low}$; let $\mathcal{G}$ be the slots in $\mathcal{I}$ with contention $\geq C_{\rm low}$ and $<C_{\rm high}$; let $\mathcal{H}$ be the slots in $\mathcal{I}$ with contention $\geq C_{\rm high}$. Combining the results from Lemmas 3.4 and 3.4, we conclude that, with high probability in $\tau$, the net delta in $\alpha_{1}N(t)+\alpha_{2}H(t)$ is at most:
$\alpha_{1}(\mathcal{A}+\mathcal{J})+O(\ln^{2}(\tau))+O(\ln^{3}(\tau))-\Omega(|\mathcal{G}|)-\Omega(|\mathcal{H}|)$
If $|\mathcal{G}|+|\mathcal{H}|>(1/10)\tau$ (and noting that $\tau>\ln^{3}(\tau)$), then the net delta in $\alpha_{1}N(t)+\alpha_{2}H(t)$ is at most $O(\mathcal{A}+\mathcal{J})-\Omega(\tau)$.
For the case that $|\mathcal{G}|+|\mathcal{H}|\leq(1/10)\tau$ (and in fact in both cases), we can omit the subtracted terms, to give an upper bound of $O(\ln^{3}\tau+\mathcal{A}+\mathcal{J})$ w.h.p. in $\tau$. ∎
5.3. Amortized behavior of $L(t)$
Consider an interval $\mathcal{I}$. Because $|\mathcal{I}|\geq(1/c_{\tau})w_{\max}(t)/\ln^{2}w_{\max}(t)$, note that $|\mathcal{I}|\geq w_{\rm min}/\ln^{2}w_{\rm min}$, where $w_{\rm min}$ is a constant of our choice that specifies the smallest window size.
We next provide bounds on how the window size of each packet changes during an interval. Proving these bounds amounts to bounding the number of times a packet listens; the actual traffic on the channel is irrelevant as the proofs are pessimistic, and hence the number of jammed slots does not appear in any of the lemma statements. Naturally, the size $\tau$ of the interval affects the number of times that a packet listens during the interval. We thus reason about a packet’s window size relative to a baseline $Z$ that satisfies $Z/\ln^{2}Z=\tau$. In some sense, $Z$ is the window size that “matches” the interval size because a packet with window $Z$ should listen $\Theta(c\ln(Z))$ times during the interval, which is the necessary number of times to change the window size by a constant factor.
The parameter $c_{\tau}$ will allow us to later tune how $Z$ relates to $w_{\max}(t)$ — in the case that $w_{\max}(t)\geq N(t)^{1/2}$, $Z$ is roughly $c_{\tau}$ times smaller than $w_{\max}(t)$. Packets with bigger windows are less likely to listen and hence less likely to change window size.
The following two lemmas captures this intuition in two ways. First, packets that start with windows smaller than $Z$ are unlikely to grow to windows that are much larger than $Z$. Second, packets with large windows (relative to $Z$) are unlikely to have their window sizes change by more than a small constant factor.
Lemma 9 (Tail bound on growth of window size).
Consider any packet during an interval $\mathcal{I}$ with $\tau=|\mathcal{I}|$. Let $Z$ satisfy $Z/\ln^{2}(Z)=\tau$. And let $W\geq w_{\rm min}$ be the initial size of the packet’s window. Let $W^{+}$ be the biggest window size the packet achieves during the interval. Then for large enough choice of constants $w_{\rm min}$ and $c$ and any $k\geq 2$: if $W\leq Z$, then
$\Pr\left[W^{+}>\Theta(kZ)\right]\leq 2^{-\Theta(c(\lg Z\cdot\lg k+\lg^{2}k))}$
or stated differently
$\Pr\left[\frac{W^{+}}{\ln^{2}(W^{+})}>\Theta(k\tau)\right]\leq 2^{-\Theta(c(\lg\tau\cdot\lg k+\lg^{2}k))}\,.$
Proof.
In order for the packet to reach window size $\Theta(kZ)$, it must first reach window size $\Theta(kZ)/2$ and then listen at least $\Theta(c\ln(kZ))$ times. So we next upper-bound the expected number of times a packet with window size at least $\Theta(kZ)/2$ listens. The expected number of times such a packet listens during the interval is
$\mu=O\left(c\cdot\frac{\ln^{3}(kZ)}{kZ}\cdot\tau\right)$ $=O\left(c\cdot\frac{\ln^{3}(kZ)}{kZ}\cdot\frac{Z}{\ln^{2}(Z)}\right)$
$=O\left(\frac{\ln^{2}(k)+\ln^{2}(Z)}{k\ln^{2}(Z)}\cdot c\ln(kZ)\right)$
$=O\left(\frac{\ln^{2}(k)}{k}\cdot c\ln(kZ)\right)$
$=O\left(\frac{1}{\sqrt{k}}\cdot c\ln(kZ)\right)$
Thus, by a standard Chernoff bound, the probability that the packet listens $\Omega(c\ln(kZ))=\Omega(\sqrt{k}\mu)$ times is upper bounded by $(1/\sqrt{k})^{\Theta(c\ln(kZ))}=(1/2)^{\Theta(c\lg(k)\cdot\ln(kZ))}$. That is,
$\Pr[W^{+}>\Theta(kZ)]\leq 2^{-\Theta(c(\lg Z\cdot\lg k+\lg^{2}k)}\ .$
Observe that $\Theta(kZ)=\Theta(k\tau\ln^{2}(Z))$ by definition of $Z$. Moreover, $\lg(Z)=\Theta(\lg\tau)$. Thus, the second statement of the bound is just multiplying both sides of the inequality $W^{+}>\Theta(k\tau\ln^{2}(Z))$ by $1/\ln^{2}(W^{+})=\Theta(1/\ln^{2}(kZ))$ when talking about reaching window size $W^{+}=\Theta(kZ)$, which cancels the $\ln^{2}(Z)$ on the right-hand side. ∎
Corollary 10 (Tail bound on growth of window sizes).
Consider an interval $\mathcal{I}$ with $\tau=|\mathcal{I}|$. Let $n$ be the number of distinct packets that are active at any point of the interval, and let $W^{+}$ denote the maximum window size reached by any packet during the interval (and hence also an upper bound on the maximum window size at the end of the interval). Then for any $k\geq 2$:
$\Pr\left[\frac{W^{+}}{\ln^{2}(W^{+})}>\Theta(k\tau)\right]\leq n\cdot 2^{-\Theta(c(\lg\tau\cdot\lg k+\lg^{2}k))}$
Proof.
This follows directly from Lemma 9 and a union bound. ∎
[…]
Proof.
Consider a packet with window size $W=\Theta(kZ)$. In order for the packet to grow or shrink its window by a factor of $e^{\gamma}$, it must listen to the channel $\Theta(\gamma\cdot c\ln(kZ))$ times. Following the same argument as Lemma 9, the expected number of times the packet listens during the interval is $\mu=O(c\ln(kZ)/\sqrt{k})$. Thus, the necessary number of listens to vary the window size by a factor of $e^{\gamma}$ is $\Omega(\sqrt{k}\gamma\mu)$, which is $\Omega(\gamma\sqrt{k})$ times the expectation. Thus, from a standard Chernoff bound, we get that the probability of this event occurring is at most $(1/(\gamma\sqrt{k}))^{\Theta(c\gamma\ln(kZ))}\leq(1/(\gamma\sqrt{k}))^{\Theta(c\gamma\ln\tau)}\leq(1/\tau)^{c\gamma\ln(\gamma k)}$, where the first step follows because $\lg\tau=\Theta(\lg Z)$. ∎
[…]
Proof.
Let $n_{0}=N(t)$ be the number of packets that are alive at the start of the interval. There are two cases:
Case 1: $\tau^{2}\geq\mathcal{A}$. By definition of the interval size, $\tau>(1/c_{\tau})\sqrt{n_{0}}$, and the total number of distinct active packets is $n=n_{0}+\mathcal{A}=O(\tau^{2})$. Thus, by Corollary 10, the probability that any packet’s window grows to $\Theta(k\tau)$ is at most $(\tau^{2})\cdot 2^{-\Theta(c(\lg\tau\cdot\lg(k)+\lg^{2}k))}=2^{-\Theta(c(\lg\tau\cdot\lg k+\lg^{2}k))}$.
Case 2: $\mathcal{A}>\tau^{2}$ and hence also $n=\mathcal{A}+n_{0}=O(\mathcal{A})$. Consider $k^{\prime}=\mathcal{A}/\tau+k$, and note that $\lg(k^{\prime})=\Theta(\lg(\mathcal{A}+k))$. Again, applying Corollary 10, the probability that any packet’s window grows to $\Theta(k^{\prime}\tau)=\Theta((\mathcal{A}/\tau+k)\cdot\tau)=\Theta(\mathcal{A}+k\tau)$ is at most $O(\mathcal{A})\cdot 2^{-\Theta(c(\lg(k^{\prime})\lg\tau+\lg^{2}(k^{\prime}))))}=2^{\Theta(\lg\mathcal{A})-\Theta(c\lg(\mathcal{A}+k)\lg\tau+\lg^{2}(\mathcal{A}+k))}\leq 2^{-\Theta(c\lg(k)\lg\tau+\lg^{2}(k))}$. ∎
[…]
Proof.
We prove this using a case analysis.
Case 1: $\Omega(\tau)$ packet injections. In this case, we argue that $L(t_{1})=O(\mathcal{A})$, which satisfies the first bullet. By Lemma 3.4, the probability that $L(t_{1})=\Omega(\mathcal{A}+2\tau)=\Omega(\mathcal{A})$ is no more than $2^{-\Theta(c\lg\tau+1)}=(1/(2\tau))^{\Theta(c)}$, where $c$ is a (large) constant specified in our algorithm.
Case 2: $\sqrt{N(t)}>L(t)$, i.e., $\tau=(1/c_{\tau})\sqrt{N(t)})$. We will show that with high probability, contention is greater than $C_{\rm low}$ in all slots of the interval, and hence the second bullet is satisfied. (In fact, contention is a fast-growing function of $N(t)$, so it is also greater than $C_{\rm high}$ as long as $N(t)$ is large enough, which can be controlled by tuning $w_{\rm min}$.) Choose any $L(t)^{2}$ packets that are active at the start of the interval. By Corollary 10, with high probability in $\tau$, none of these packets ever reach window size above $O(\tau\log^{2}(\tau))=O(N(t)^{1/2}\log^{2}(N(t)))$. Moreover, at most $\tau=O(\sqrt{N(t)})$ of the packets can succeed during the interval. Thus, the contention in every slot is at least $\Omega(N(t)^{1/2}/\log^{2}(N(t)))\gg C_{\rm low}$.
Case 3: $\mathcal{A}=O(\tau)$ and $N(t)^{1/2}\leq L(t)$, i.e., $\tau=(1/c_{\tau})\cdot{\frac{w_{\max}(t)}{\ln^{2}(w_{\max}(t))}}$. We will argue that either the first or second bullet holds.
We begin by arguing that packets with small windows initially are irrelevant. Observe that the total number of packets active during the interval is at most $O(N(t)+\mathcal{A})=O(\tau^{2})$. Thus, for $c_{\tau}$ sufficiently large [^10] Choosing large $c_{\tau}$ means that “$Z$” in the preceding lemmas is smaller relative to $w_{\max}(t)$. Specifically, $w_{\max}(t)$ is roughly $c_{\tau}$ times larger than $Z$. In other words, a window of size close to $w_{\max}(t)$ corresponds to using $k$ close to $c_{\tau}$ in the lemmas., we can apply Corollary 10 and Lemma 3.4 to get high probability bounds. In particular, with failure probability at most $1/\tau^{\Theta(1)}$, no packet with window smaller than say $(3/4)w_{\max}(t)$ grows to window size larger than say $(7/8)w_{\max}(t)$. (The particular choice of constants (3/4) and (7/8) is not important — any constant can be achieved by tuning $\gamma$ in Lemma 3.4; e.g., $\gamma=1/7$ for the stated constants. [^11] Tuning $\gamma$ also places restrictions on $c$ and $c_{\tau}$ as we would like $c,c_{\tau}\gg 1/\gamma$ and hence $c,k\gg 1/\gamma$ for the failure probability in Lemma 3.4 to be polynomially small in $\tau$.)
For the remainder, we focus on any packet that starts the interval with window size at least $(3/4)w_{\max}(t)$. By Lemma 3.4, with high probability such a packet maintains a windows size between say $(1/2)w_{\max}(t)$ and $(6/5)w_{\max}(t)$, meaning that the probability of listening does not vary by more than a constant factor. In particular, the packet chooses to listen to each slot with probability at least $(1/2)c\ln^{3}(w_{\max}(t))/w_{\max}(t)$ and at most $(6/5)c\ln^{3}(w_{\max}(t))/w_{\max}(t)$
Now suppose that at least a (9/10)-fraction of the slots in $\mathcal{I}$ have low contention and are not jammed. We next count two things: $x_{s}$ is the number of times the packet listens and hears silence, and $x_{n}$ is the number of times the packet listens and hears noise. As long as $x_{s}-x_{n}\geq\delta(c\ln(w_{\max}(t)))$, for constant $\delta$, then the packet’s window shrinks by a factor of roughly $e^{\delta}$, which corresponds to the constant $d$ in the first bullet of the lemma statement. We thus need only argue that $x_{s}-x_{n}=\Omega(c\ln(w_{\max}(t)))$ with high probability (and then take a union bound across packets) to justify that such a constant exists.
There is one subtle detail here: the packet’s window has to be large enough that the lower bound we derive on $x_{s}$ below would not result in the packet’s window dropping below the minimum window size $w_{\rm min}$; it would be sufficient for the following if $w_{\max}(t)\geq 2w_{\rm min}$, for example. We can enforce this constraint by simply setting $C_{\rm low}<1(/2w_{\rm min})$, as even a single packet with window below $2w_{\rm min}$ would contribute more than $C_{\rm low}$ to the contention.
We first argue that the packet is likely to observe many empty slots. By Lemma 2 and the fact that $C_{\rm low}\leq 1/w_{\rm min}\leq 1/20$, each of the unjammed low-contention slots is empty with probability at least $9/10$. Thus, the expected value of $x_{s}$ is at least $(9/10)\cdot(1/2)c\ln^{3}(w_{\max}(t))/w_{\max}(t)\cdot(9/10)(1/c_{\tau})w_{\max}(t)/\ln^{2}(w_{\max}(t))=(81/200)\cdot(c/c_{\tau})\ln(w_{\max}(t))$. Now we can apply an adaptive Chernoff bound to conclude that $x_{s}\geq(1/3)\cdot(c/c_{\tau})\ln(w_{\max}(t))$ with high probability in $w_{\max}(t)\geq\tau$.
We next consider the number noisy slots, which has two components. First, for slots where the contention exceeds $C_{\rm low}$ or the slot is jammed, the worst case is that all such slots are full. The expected number of listens to such slots is at most $(6/5)c\ln^{3}(w_{\max}(t))/w_{\max}(t)\cdot(1/10)(1/c_{\tau})w_{\max}(t)/\ln^{2}(w_{\max}(t))=(6/50)\cdot(c/c_{\tau})\ln(w_{\max}(t))$, which is at most $(1/8)(c/c_{\tau})\ln(w_{\max}(t))$ with high probability. There are also the slots where the contention is low but that happen to be noisy. For low contention unjammed slots, there may also be noise. But the probability of noise is at most $(1/10)$. The expected number of noisy unjammed low-contention slots that the packet listens to is at most $(1/10)\cdot(6/5)c\ln^{3}(w_{\max}(t))/w_{\max}(t)\cdot(1/c_{\tau})w_{\max}(t)/\ln^{2}(w_{\max}(t))=(3/25)\cdot(c/c_{\tau})\ln(w_{\max}(t))$. Thus, by a Chernoff bound, with high probability, the number of such slots is at most $(1/8)(c/c_{\tau})\ln(w_{\max}(t))$. Adding these two together, we get that $x_{n}\leq(1/4)\cdot(c/c_{\tau})\ln(w_{\max}(t))$, with high probability.
In conclusion, with high probability, $x_{s}-x_{n}\geq(1/3-1/4)\cdot(c/c_{\tau})\ln(w_{\max}(t))=(1/(12c_{\tau}))\cdot c\ln(w_{\max}(t))$. ∎
5.4. Bounding the decrease in $\Phi(t)$ over intervals by combining $N(t)$, $H(t)$, and $L(t)$
We now consider the change to $\Phi$ over an interval by combining the bounds for $N$, $H$, and $L$. The first lemma here states that with high probability, if there are not too many packet arrivals or jammed slots, then $\Phi$ is very likely to decrease by an amount that is proportional to the size of the interval. There is a low probability failure event in which the desired decrease to potential does not occur. The second lemma focuses on such failures, providing a tail bound for the likelihood that $\Phi$ increases by a large amount. Together, these lemmas match the conditions necessary for the betting game.
[…]
Proof.
First, consider the high probability bounds on how much the terms in the potential increases during $\mathcal{I}$.
  • In the worst case, $N(t)$ increases by at most $\mathcal{A}$.
  • By Lemma 3.4, $\alpha_{1}N(t)+\alpha_{2}H(t)$ increases by at most $O(\ln^{3}\tau+\mathcal{A}+\mathcal{J})$, with high probability in $\tau$.
  • By Lemma 9, $L(t)$ increases by at most $O(\mathcal{A}+\tau)$, with high probability in $\tau$.
We next consider two cases to conclude that the potential decreases overall.
Case 1: at least $(9/10)$ fraction of the slots have low contention. Then Lemma 3.4 states that $L(t)$ decreases by $\Omega(\tau)-O(\mathcal{A}+\mathcal{J})$. Adding the increases summarized above to the other terms, we get a net decrease of $\Omega(\tau)-O(\mathcal{A}+\mathcal{J}+\ln^{3}\tau)=\Omega(\tau)-O(\mathcal{A}+\mathcal{J})$.
Case 2: at most a $(9/10)$ fraction of the slots have low contention. Then Lemma 3.4 states that $\alpha_{1}N(t)+\alpha_{2}H(t)$ decreases by at least $\Omega(\tau)-O(\mathcal{A}+\mathcal{J})$. Choosing $\alpha_{1}$ and $\alpha_{2}$ much larger than $\alpha_{3}$ the $\Omega(\tau)$ decrease to the first two terms dominates the $O(\tau)$ increase to the third term. ∎
Our next lemma provides a tail bounds on the amount by which $\Phi(t)$ increases, which is relevant if the high probability bound in the preceding lemma fails. At first glance, the reader may be surprised to see that the amount of jamming does not occur in the lemma statement. This omission is in part due to the fact that any jamming is subsumed by the $\Theta(k\tau^{2})$ term because there can be at most $\tau$ jammed slots.
[…]
Proof.
In fact, we will upper-bound $\Phi(t^{\prime})$, which implies an upper bound on $\Phi(t^{\prime})-\Phi(t)$ because $\Phi$ is always nonnegative. We begin by providing worst-case upper bounds on $N(t^{\prime})$ and $H(t^{\prime})$. We complete the proof by incorporating the tail bound on $L(t^{\prime})$.
Observe that $N(t^{\prime})\leq N(t)+\mathcal{A}$. Because $\tau\geq(1/c_{\tau})\sqrt{N(t)}$, we have $N(t^{\prime})\leq(c_{\tau}\tau)^{2}+\mathcal{A}=O(\tau^{2})+\mathcal{A}$. To bound $H$, notice that each packet active at time $t^{\prime}$ contributes at most $11/\ln(w_{\rm min})<$ 1 to $H(t^{\prime})$. Thus, $H(t^{\prime})<N(t^{\prime})=O(\tau^{2})+\mathcal{A}$.
Finally, by Lemma 3.4, the probability that $L(t^{\prime})$ exceeds $\Theta(\mathcal{A}+k\tau)$ is upper bounded by $2^{-\Theta(c(\lg\tau\cdot\lg k+\lg^{2}k))}$. And when $L(t^{\prime})$ falls below the stated threshold, we have $\Phi(t^{\prime})\leq\alpha_{1}\cdot O(\mathcal{A}+\tau^{2})+\alpha_{2}\cdot O(\mathcal{A}+\tau^{2})+\alpha_{3}\cdot O(\mathcal{A}+k\tau)=O(\mathcal{A}+\tau^{2}+k\tau)=O(\mathcal{A}+k\tau^{2})$. ∎
5.5. Using the analysis of $\Phi(t)$ to prove throughput via a betting-game argument
The preceding section establishes progress guarantees over sufficiently large intervals in the form of Theorems 3.3 and 3.4. Specifically, consider an interval $\mathcal{I}$ starting at slot $t$ with length $|\mathcal{I}|=\tau=(1/c_{\tau})\cdot\max\big{\{}\frac{w_{\max}(t)}{\ln^{2}(w_{\max}(t))},N(t)^{1/2}\big{\}}$. Then, with high probability in $|\mathcal{I}|$, $\Phi(t)$ decreases over $\mathcal{I}$ by at least $\Omega(\tau)-O(\mathcal{A}+\mathcal{J})$, where $\mathcal{A}$ is the the number of packet arrivals and $\mathcal{J}$ is the number of jammed slots in $\mathcal{I}$. Critically, these bounds hold with high probability in $\tau$. In this section we show how to apply these lemmas to achieve results for the execution with high probability in $N$, the total number of packets.
Since the adversary is adaptive, we have to be careful in combining bounds across intervals. The adversary can use the results of earlier intervals in choosing new arrivals and jamming, which affects the size of later intervals. This may lead to a more advantageous interval length for the adversary, which might increase its probability of success.
To reason about this process, we reframe it in a setting that resembles a random walk, which we describe below in a betting game. Our analysis of this game then allows us to analyze the implicit throughput (recall Section 1.1).
The Betting Game. We first summarize the betting game and then later relate it to the backoff process. The adversary corresponds to a bettor who makes a series of bets. Each bet has a size equal to the duration $\tau$. The bettor also has some amount of money, which is initially 0 dollars. When the bettor loses a bet, the bettor loses some money, and when the bettor wins, the bettor wins some money. (The amounts won or lost are specified below as a function of the size of the bet.) Additionally, at any time, the bettor may choose to receive a passive income. The passive income is added to the bettor’s wealth. The total amount of passive income taken, however, means that the bettor must play the game longer. The game begins when the bettor first takes some passive income, and the game does not end until either the bettor goes broke or the bettor has resolved bets totaling $\Omega(P)$ size, where $P$ is the passive income received, whichever comes first. The bettor’s goal is to complete the game without going broke. Importantly, although the bettor can always choose to take more passive income, doing so increases the total play time.
We set the details of the betting game to mirror the backoff process. Each bet corresponds to an interval. Passive income during a bet corresponds to the number of arrivals and jammed slots during the interval. Money corresponds to potential.
The bettor loses a size-$\tau$ bet with probability at least $1-\frac{1}{\text{poly}(\tau)}$. If the better loses the size-$\tau$ bet, it loses $\Theta(\tau)$ dollars. This loss corresponds to the high-probability event (in $\tau$) of Theorem 3.3. The bettor wins a size-$\tau$ bet with probability $O(1/\text{poly}(\tau))$. If the bettor wins the bet, it gets $\Theta(\tau^{2})$ dollars, plus $Y$ bonus dollars, where $Y$ is a random variable such that $\Pr[Y\geq k\tau^{2}]\leq\frac{1}{\text{poly}(\tau)}\cdot 2^{-\Theta(\log^{2}k)}$; these winnings correspond to tail bound of Theorem 3.4. (Of course, during each bet, the bettor can may also gain passive income for arrivals and jammed slots.)
We pessimistically give the bettor the power to choose arbitrary bet sizes (subject to a minimum interval size, which itself is determined by $w_{\rm min}$), and the bettor is even allowed to place bets whose loss would cause the bettor to end with negative money. (In the actual backoff process, the interval sizes are dictated by the current state of the system, and not entirely under the control of the adversary.)
The rules of betting game are set pessimistically (in favor of the bettor) such that when the bettor wins, $\Phi(t)$ increases more slowly than the bettor’s wealth increases, and when the bettor loses, $\Phi(t)$ decreases at least as fast as the bettor’s wealth decreases. Therefore, this betting game stochastically dominates the potential function.
The takeaway is that at any point $t$, the bettor’s wealth is an upper bound on $\Phi(t)$. Because $\Phi(t)$ is an upper bound on the number of packets in the system, the bettor going broke corresponds to all packets succeeding. We thus obtain good implicit throughput, because there must either be many jammed slots or packet arrivals, or there must be many packets succeeding, leading to inactive slots.
Upper bounding the bettor’s maximum wealth/potential and showing $\Omega(1)$ implicit throughput.
In Lemma 3.4, we provide a high-probability upper bound on the bettor’s maximum wealth and the amount of time until they go broke, which corresponds to there being no packets in the system. We use this to establish Corollary 11, which proves $\Omega(1)$ implicit throughput. Finally, Corollary 12 bounds the potential function in terms of the number of packet arrivals and jammed slots (which, in terms of the betting game, correspond to passive income).
[…]
Proof.
Group the bets into bet classes, where bet class $j$ contains all bets with size in $[2^{j},2^{j+1})$.
Passive income. For the purpose of determining how much passive income the bettor takes (and hence how long to play the game), we will let the adversary have a limited view of the future: for each bet class, the bettor can see the sequence of outcomes for bets in that bet class; with this information the bettor can plan how many bets to allocate to each class and how much passive income to take.
Let $P$ be the total passive income that the adversary chooses to take as a result of this information; we will generously allow the adversary to take that passive income at the very beginning of the game.
We will show w.h.p. in $P$, that the bettor goes broke over the first $O(P)$ slots, and never has more than $O(P)$ dollars.
Large bets. Before proceeding further, we will dispose of the large bets. A bet of size $P^{1/4}$ or larger is big, and otherwise the bet is small. With high probability in $P$, the bettor loses every big bet: If the bet size is at least $P^{1/4}$, then the probability of winning the bet is $O(1/\text{poly}(P^{1/4}))$ (by the rules of the betting game.) Since there can only be $\text{poly}(P)$ bets of this size before going broke, by a union bound, none of them are successes w.h.p., for appropriate choice of polynomials.
Analyzing a bet class. We now consider the bets in a single bet class. We will imagine that the bets in that bet class are an infinite stream of possible bets, and that the adversary can see the outcomes in advance when choosing whether or not to make the next bet. (This is to account for the adaptive nature of the adversary that it might choose to make bets of a different size based on past history.) We will show that every prefix of this stream of bets is “good” with high probability.
Bonus dollars. First, we will examine the bonus dollars. We will show that, with high probability, for any prefix of the stream of bets: if the bettor wins $X$ regular dollars, then it also wins $O(X)$ bonus dollars, with high probability in $P$. Recall that for any single bet in bet class $[2^{j},2^{j+1})$, the bonus dollars $Y$ are drawn from a distribution where $\Pr[Y>k\cdot 2^{j+1}]\leq 1/2^{\Theta(j+\ln^{2}k)}$, by the rules of the game.
First, we consider the first $O(\log(P))$ successful bets in each bet class of size at most $P^{1/4}$. The probability of winning at least $\Theta(P^{1/2})$ bonus dollars is polynomially small in $P$, and there are at most $O(\log(P))$ such winning bets. So we conclude that the bonus dollars from all such bets is at most $O(P)$ with high probability in $P$.
Next, we fix a specific bet class $j$ and consider some fixed prefix of the bet sequence for this bet class. Let $B$ be the set of bets that the bettor wins; we assume $B>\Theta(\log{P})$, as we have already accounted above for the first few successful bets. We want to examine each possible value of $k$ for the bettor’s “bonus multiplier”, hence for each $k\in\mathbb{Z}^{+}$ define the following indicator random variables: for the $i$th bet in $B$, let $x_{i}^{(k)}=1$ iff the $i$th bet in $B$ yields at least $k2^{j+1}$ bonus dollars. (Otherwise, $x_{i}^{(k)}=0$.)
The total number of bonus dollars won from bets within this bet class is at most $\sum_{i=1}^{|B|}\sum_{k=1}^{\infty}2^{j+1}x_{i}^{(k)}$. As we only want to consider values of $k$ that are a power of 2, we can round up the bonus dollar winnings and conclude that the bonus dollars are bounded by: $2^{j+2}\sum_{k\in\{2^{i}\mid i\in\mathbb{Z}_{\geq 0}\}}\sum_{i=1}^{|B|}kx_{i}^{(k)}$.
Given the probability of winning bonus dollars, we know that ${\rm E}\left[x_{i}^{(k)}\right]\leq 2^{-\Theta(j+\log^{2}{k})}\leq 2^{-\Theta(\log^{2}{k})}$. We conclude that ${\rm E}\left[k\sum_{i=1}^{|B|}kx_{i}^{(k)}\right]\leq k|B|2^{-\Theta(\log^{2}{k})}\leq|B|2^{-\Theta(\log^{2}{k})}$.
Thus, the expected total bonus dollars for the bet class are:
$2^{j+2}\sum_{k\in\{2^{i}\mid i\in\mathbb{Z}_{\geq 0}\}}|B|2^{-\Theta(\log^{2}{k})}=O(2^{j+2}|B|)$
By Theorem 4, we conclude that the total bonus dollars for the bet class is $O(2^{j+2}|B|)$ probability at least $1-e^{-\Theta(|B|)}$, where $B>\Theta(\log P)$. Since this decreases exponentially as $B$ increases, it holds for all prefixes of the bet sequence.
Since $|B|$ bets in bet class $j$ make at least $\Theta(2^{j}|B|)$, and with high probability in $P$ there are at most $O(2^{j}|B|)$ bonus dollars, we conclude that with high probability we can bound the total bonus dollars in terms of: (i) $O(P)$ bonus dollars (for the first few bets), plus (ii) $O(X)$ bonus dollars for a bet class that that wins $X$ dollars.
Wins versus losses. Now we analyze the total winnings over any prefix of the bets in a given bet class.
The first observation is that the first $O(\log(P))$ bets result in at most $O(\log(P))$ wins, and hence at most $O(P^{1/4}\log(P))=O(P)$ income. We will see that the bettor never makes more money than this.
Next, we will show that once $\Omega(\log(P))$ bets have been made, w.h.p. in $P$: at all subsequent points in the betting sequence (for that bet class), all but at most an $\alpha$-fraction of bets are losses, where $\alpha$ is a constant that we can choose by changing the polynomial.
More precisely, recall that the bettor wins a bet in this bet class with probability 1/poly($21/\text{poly}(2^{j})\leq\alpha/$ 2. Thus in the prefix containing $B$ bets, the expected numbers of wins by the bettors is at most $B\alpha/2$. We conclude by a straightforward application of Theorem 4 that the bettor wins more than $\alpha B$ bets with probability at most $e^{-\Theta(B)}$. Since this is exponentially decreasing in $B$, it holds for every prefixes of bets in the bet class that is longer than $\Omega(\log(P))$.
Imagine, for example, that the bettor makes $B$ bets in bet class of sizes $[2^{j},2^{j+1})$. If $B=\Omega(\log(P))$, then we conclude that (with high probability in $P$) the bettor loses $\Theta(2^{j})(1-\alpha)B-\Theta(2^{j+1})\alpha B$—remembering that we have already shown that if the bettor wins $X$ regular dollars in this prefix, then it also wins only $O(X)$ bonus dollars. Thus for a proper choice of $\alpha$, the bettor loses money for every choice of $B$.
Specifically, if the better chooses to bet $X$ in a small bet class and those bets consists of more than $O(\log(P))$ bets, then with high probability the bettor loses $\Theta(X)$ dollars.
Wrapping up the proof. Thus we conclude that after bets totalling $\Theta(P)$ either $\Theta(P)$ of the bets are large—and the bettor goes bankrupt—or $\Theta(P)$ of the bets are small. In the latter case, at least $\Theta(P)$ of those bets must not consist of “the first $O(\log(P))$ bets in a bet class”, since there are only $\log^{2}(P)$ such bets. Thus with high probability the bettor loses $\Theta(P)$ dollars and goes bankrupt. ∎
Corollary 11 (Constant implicit throughput).
At the $t$-th active slot in the betting game, with high probability in $t$, there has been a passive income of $\Omega(t)$ dollars. Hence, the implicit throughput is $\Omega(1)$.
Proof.
We briefly explain here how Lemma 3.4 implies implicit throughput. Consider a time horizon $t$, and suppose that the bettor has received $P=t/c$ dollars from passive income, for constant $c$ matching the big-$O$ of the lemma. Then from Lemma 3.4, with high probability in $t/c$, the bettor goes broke within $c\cdot t/c=t$ time; that is, there are no active packets at time $t$. We thus obtain the $\Omega(1)$ throughput result. ∎
We conclude by noting that the above corollary yields Theorem 3. Moreover, Corollary 4 is a special case of Theorem 3.
Our next corollary just restates part of Lemma 3.4 with respect to the potential. Due to Theorems 3.3 and 3.4, the potential increases less than the bettor’s net worth does, and it decreases at least as fast. That is, the betting game stochastically dominates the potential.
Corollary 12 (Bound on potential in terms of packet arrivals and jammed slots).
Suppose that we have had $N_{t}$ packet arrivals and $\mathcal{J}_{t}$ jammed slots up until slot $t$. Then w.h.p. in $N_{t}+\mathcal{J}_{t}$, for all slots $t^{\prime}=1,\ldots,t$, the potential $\Phi(t^{\prime})=O(N_{t}+\mathcal{J}_{t})$.
Bounds on throughput for infinite and finite streams.
We next address the case of adversarial queuing arrivals with granularity $S$, and where the arrival rate $\lambda$ is a sufficiently small constant. Recall that the “arrival rate” $\lambda$ limits the number of packet arrivals and jammed slots to at most $\lambda S$ in every interval of $S$ consecutive time slots.
We start with Lemma 13, showing that, in the case of adversarial queuing arrivals with a sufficiently small arrival rate, there is an inactive slot within any interval of length $S$. This is used to prove Corollary 14, a high-probability bound on the number of packets in the system under an adversarial queuing arrival model, which is the same as Corollary 5.
Lemma 13 (Existence of inactive slot).
Consider an input stream of adversarial queuing arrivals with granularity $S$ and sufficiently small arrival rate $\lambda=O(1)$. Consider also any particular time slot $t$. Then with high probability in $S$:
  • there is an inactive slot at some time $t^{\prime}$ with $t-\Theta(S)\leq t^{\prime}<t$, and
  • there is an inactive slot at some time $t^{\prime}$ with $t<t^{\prime}\leq t+\Theta(S)$.
Proof.
Consider a time $t$, and let $t^{\prime}<t$ be the last inactive slot before time $t$. Our goal is to prove that $t-t^{\prime}=O(S)$, with high probability in $S$
Let us first consider the implication of large $t-t^{\prime}\geq S$ on implicit throughput. Then we have a time period where at most $\lambda(t-t^{\prime}+S)\leq 2\lambda(t-t^{\prime})$ packet arrivals or jamming occur. Since there are packets active throughout this entire interval from $t^{\prime}$ to $t$ by construction, the implicit throughput is at most $2\lambda(t-t^{\prime})/(t-t^{\prime})=O(\lambda)$ at time $t$. If $\lambda$ is small enough, this level of throughput does not meet the $\Omega(1)$ implicit throughput promised by Corollary 11, which contradicts the assumption that $t-t^{\prime}$ is large. We can thus apply Corollary 11 at time $t$ to conclude that with high probability in $S$, the $\Omega(1)$ implicit throughput is met at $t$, and hence $t-t^{\prime}=O(S)$.
To achieve the second claim, we apply the same argument at time $t+\Theta(S)$, which implies an inactive slot between $t$ and $t+\Theta(S)$, w.h.p. ∎
We can leverage the first bullet point of Lemma 13 to obtain the following bound on an adversary with granularity $S$. (And we use the second bullet later in Theorem 16; although it immediately follows from the first bullet point, it is still useful to note explicitly.)
Corollary 14 (Bounded backlog for adversarial-queuing arrivals).
Consider an input stream of adversarial queuing arrivals with granularity $S$ and sufficiently small arrival rate $\lambda=O(1)$. Consider any particular slot $t$. Then with high probability in $S$, the number of packets in the system at time $t$ is $O(S)$.
Proof.
By Lemma 13, there is an inactive slot at time $t-O(S)$ with high probability in $S$. If that is the case, there can be at most $O(S)$ packet arrivals plus jammed slots since the last inactive slot. ∎
This completes the proof Corollary 5. Again, since Corollary 4 is a special case of Theorem 3, we have proved our central results on throughput.
5.6. Channel access/energy bounds
Energy bounds for finite streams.
Our first theorems apply for finite instances.
[…]
Proof.
From Corollary 4, the number of active slots is $O(N+\mathcal{J})$. We also know from Corollary 12 that the potential function is $O(N+\mathcal{J})$ throughout the entire execution w.h.p. in $N+\mathcal{J}$.
One consequence is that for all $t$, we can bound the largest window size ever reached, denoted by $W^{*}$. In particular, at each time $w_{\max}(t)=O(\text{poly}(N+\mathcal{J}))$ w.h.p., and hence $W^{*}=O(\text{poly}(N+\mathcal{J}))$ w.h.p. (In fact, for all $t$, $w_{\max}(t)=O((N+\mathcal{J})\log^{2}(N+\mathcal{J}))$ w.h.p., but we do not need to be that exact.)
First, let us count the number of times a packet can back off before it reaches a window size of $W^{*}$, assuming no backons. Suppose a packet has window size $w$. Then each time it backs off, its window increases in size by factor of $(1+1/\Theta(\log w))$. Thus, there are $\Theta(\log w)$ backoffs before the window size doubles. Rounding $w$ up to the maximum $W^{*}$, we get $O(\log(W^{*}))$ backoffs to double the window size and $O(\log^{2}(W^{*}))$ backoffs to reach maximum window size.
In order for a packet to perform additional backoffs, it must also back on as otherwise its window would increase beyond $W^{*}$. In particular, for each additional backoff, the packet must also perform $\Omega(1)$ backons.
Next, we shall bound $k$, the number of backons that the packet performs, with high probability. Since each backon allows for $O(1)$ additional backoffs, the total number of backoffs is now given by $O(\log^{2}(W^{*})+k)$. And the number of listening attempts is the number of backoffs plus backons, and also $O(\log^{2}(W^{*})+k)$.
The key point is that each time a backon occurs, the slot is empty. So the only question is how many times the packet can listen to empty slots before it chooses to also sends. Every time a packet listens, there is an $\Omega(1/\log^{3}W^{*})$ probability that it also makes a sending attempt. The probability that the packet never sends on these empty slots is $(1-\Omega(1/\log^{3}W^{*}))^{k}$, which is polynomially small in $W^{*}$ for $k=\Theta(\log^{4}W^{*})$. Thus with high probability in $W^{*}$, we have $*k=O(\log^{4}W^{*}$).
To conclude the proof, we substitute in that $W^{*}=O(\text{poly}(N+\mathcal{J}))$ w.h.p. ∎
Theorem 15 (Energy bounds for the finite case against an adaptive and reactive adversary).
Consider an input stream with $N$ packets and $\mathcal{J}$ jammed slots. Assume that the adversary is adaptive and reactive. Then, w.h.p. in $N+\mathcal{J}$
  • a packet accesses the channel at most $O((\mathcal{J}+1)\log^{3}(N+\mathcal{J})+\log^{4}(N+\mathcal{J}))$ times.
  • the average number of channel accesses is $O((\mathcal{J}/N+1)\log^{4}(N+\mathcal{J}))$ times.
Proof.
We first explain the worst-case bound. Recall that with an adaptive adversary, the jammer commits at the start of a slot whether that slot will be jammed. In contrast, the reactive adversary does not decide to jam a slot until after it sees which packets have chosen to send during that slot. That is, the reactive adversary may reserve all of its jamming for slots when a packet would have otherwise successfully transmitted. Recall also that the reactive adversary cannot react to listening—it can only react to sending.
We revisit the proof of Theorem 3.4 at the point that reactivity has an impact—bounding $k$, the number of backons, before the packet successfully transmits. Each time a packet observes an empty slot, it also chooses to send with probability $\Omega(1/\log^{3}W^{*})$. Notice that if the packet observes $\rho\log^{3}W^{*}$ empty slots, then it would choose to send in $\Omega(\rho)$ of them in expectation. But the adversary can only block $\mathcal{J}$ of these, so choosing $\rho=\Omega(\mathcal{J})$ is large enough to get more than $\mathcal{J}$ sends, and hence a send that is not jammed, in expectation. To achieve a high probability result, we apply a Chernoff bound: with failure probability exponentially small in $\rho$, the packet chooses to send $\Omega(\rho)$ times. Choosing $\rho=\Theta(\mathcal{J}+\log W^{*})$ for some sufficiently large constant, we get that with high probability $W^{*}$ the number of sends is more than $\mathcal{J}$. Thus, with high probability in $W^{*}$, $k\leq\rho\log^{3}W^{*}=O((\mathcal{J}+\log W^{*})\log^{3}W^{*})$.
It is an additional step to compute the average bound. Now, collectively, there must be $N$ successful transmissions. But since the jammer can block up to $\mathcal{J}$ transmissions that otherwise would be successful, there need to be enough listening attempts so that there would be $N+\mathcal{J}$ successful sends without the jammer. The bounds follow immediately by a similar argument. ∎
Infinite streams: with adversarial queueing and in general.
We first turn to infinite streams with adversarial-queueing arrivals. The main tool is Lemma 13, which allows us to transform the adversarial queuing case into finite instances that are not very large.
Theorem 16 (Energy bound for adversarial-queuing arrivals against an adaptive adversary).
Consider an input stream of adversarial queuing arrivals with granularity $S$ and sufficiently small arrival rate $\lambda=O(1)$. Consider an adaptive but not reactive adversary. Then w.h.p. in $S$, a particular packet accesses the channel at most $O(\log^{4}S)$ times.
Proof.
Consider a particular packet $p$ injected at time $t$. Now let us consider the latest timestep $t_{0}<t$ with no active packets and the earliest timestep $t_{1}>t$ with no active packets (and hence packet $p$ has also finished). We thus have a finite instance with $N+\mathcal{J}\leq t_{1}-t_{0}$. By Lemma 13, $t_{1}-t_{0}=O(S)$, with high probability in $S$. The current claim then follows from Theorem 3.4. ∎
We next consider an adversary that is both adaptive and reactive for adversarial queuing. The argument is similar, except that now we apply the finite case of Theorem 15.
Theorem 17 (Energy bounds for adversarial-queuing arrivals against an adaptive and reactive adversary).
Consider an input stream of adversarial-queuing arrivals with granularity $S$ and sufficiently small arrival rate $\lambda=O(1)$. Consider a reactive and adaptive adversary.
  • W.h.p. in $S$ a packet accesses the channel at most $O(S)$ times.
  • W.h.p. in $S$, the average number of channel accesses per slot is $O(\log^{4}S)$, averaging over all the slots.
Proof.
For the first claim, we leverage the second part of Lemma 13. That is, consider a packet injected at time $t$. Then, with high probability in $S$, there is an inactive slot before time $t+\Theta(S)$, and hence the packet finishes by time $t+O(S)$. Even if it accesses the channel in every step during which it is active, the packet only makes $O(S)$ accesses.
For the second claim, we reduce to the finite case, but the way we do this is slightly different as we want to be able to average the total number of accesses across a reasonably large interval. Consider times $t_{i}=i\cdot\Theta(S)$, and consider a particular interval from time $t_{i}$ to time $t_{i+1}$. Let $t_{i}^{\prime}<t_{i}$ be the latest inactive slot before $t_{i}$ and $t_{i+1}^{\prime}>t_{i+1}$ be the earliest inactive slot after $t_{i+1}$. Then we consider the interval from $t_{i}^{\prime}$ to $t_{i+1}^{\prime}$ as a finite instance. By Lemma 13, the total size of this interval is $\Theta(S)$ with high probability in $S$. There are thus $O(S)$ packets and $O(S)$ jamming. We can then apply Theorem 15 to this finite instance to conclude that the total number of channel accesses is thus at most $O(S\log^{4}(S))$. Even if all of these accesses fall in the subinterval from $t_{i}$ to $t_{i+1}$, each slot in that subinterval receives $O(\log^{4}(S))$ accesses on average. ∎
[…]
Proof.
Since we are only looking at events occurring before time $t$, we essentially have a finite instance. That is, consider an extension to the execution after time $t$ where the adversary performs no further jams or injections. Then Theorem 3.4 and Theorem 15 apply, and they achieve the bounds claimed here except across the entire extended execution. Truncating the execution at time $t$ only reduces the number of channel accesses. ∎
6. More Related Work
Contention resolution has been an active area of research for several decades. While we have already discussed closely-related work in Sections 1, here we expand on our discussion and also summarize additional related results from the literature.
Robustness. Early work on jamming resistance by Awerbuch et al. (awerbuch:jamming, ) considers the case of $N$ devices, each of which always has a packet to ready to send; that is, rather than changing over time, the number of packets contending for channel access is always $N$. For any interval of at least $T$ consecutive slots, at most a $(1-\lambda)$-fraction may be jammed, where the constant $\lambda>0$ is unknown. In this setting, Awerbuch et al. (awerbuch:jamming, ) showed that constant throughput is possible with a polylogarithmic number of transmissions per packet. Several subsequent results have built on this result, addressing multi-hop (richa:jamming2, ; richa:competitive-j, ) and co-located networks (richa:jamming4, ), reactive jamming (richa:jamming3, ; richa:efficient-j, ), and aspects of signal propagation (ogierman:competitive, ). However, this series of results relies on devices knowing rough bounds on $N$ and $T$; in particular, devices must set a parameter whose value is $O(1/(\log\log(N)+\log(T)))$.
A similar restriction on jamming is explored by Anantharamu et al. (anantharamu2019packet, ; anantharamu2011medium, ), where $\rho t+b$ slots in any window of consecutive $t$ slots can be jammed, for $0<\rho\leq 1$, and a non-negative integer, $b$. Probabilistic jamming has also been considered by Chlebus et al. (Chlebus:2016:SWM:2882263.2882514, ) in a multi-channel setting, and by Agrawal et al. (DBLP:conf/spaa/AgrawalBFGY20, ) where packets have delivery deadlines. When the amount of jamming can be unbounded, Bender et al. (BenderFiGi19, ; BenderFiGi16, ) showed expected constant throughput is possible over non-jammed slots using a $\operatorname{polylog}(N)$ number of sending attempts per packet. Chang et al. (ChangJiPe19, ) give an elegant algorithm that achieves the same asymptotic result, but drastically improves the expected throughput, rivaling the $1/e$ result achieved by the well-known slotted ALOHA (binder:aloha, ), despite the more-challenging adversarial setting.
There is a growing body of work (ChenJiZh21, ; jiang:robust, ; DeMarcoSt17, ; BenderKoKu20, ) on how to design contention resolution protocols where packets listening on the channel during a given slot learn whether or not that slot was successful, but not whether the slot was empty or noisy; that is, packets lack the ability to detect collisions. As mentioned earlier, the results by De Marco and Stachowiak (DeMarcoSt17, ), De Marco, Kowalski, and Stachowiak (marco:time, ), and Bender et al. (BenderKoKu20, ) address this setting. Dealing with jamming in the absence of collision detection is challenging. Here, results by Chen et al. (ChenJiZh21, ) and Jiang and Zheng (jiang:robust, ) show how jamming can still be tolerated, although throughput is $\Theta(1/\log N)$ and packets are allowed to listen on the channel for free.
There is a large body of work on applied security considerations with respect to jamming. The impact of these attacks has been evaluated under a variety of wireless settings and adversarial strategies (bayraktaroglu:performance, ; xu:feasibility, ; lin:link, ; aschenbruck:simulative, ), and several defenses have been proposed. For example, spread-spectrum techniques involve devices coordinating their communication over different portions of the spectrum in order to evade collisions (spread:foiling, ; navda:using, ; deejam:wood, ). Another defense is to have devices map the areas impacted by jamming and then routing traffic around those areas (wood2003jam, ; han2018hierarchical, ). Specialized antenna technology can be used to filter out a jamming signal, which is a process referred to as null steering (zheng2018design, ; mcmilin2015gps, ).
Energy efficiency. As discussed in Section 1, there are a handful of results on contention resolution that achieve constant throughput and address full energy efficiency. To reiterate, Bender et al. (BenderKoPe18, ) achieve expected $O(\log(\log^{*}N))$ channel accesses per packet, which Chang et al. (ChangKoPe19, ) prove is tight. De Marco and Stachowiak (DeMarcoSt17, ) and De Marco, Kowalski, and Stachowiak (marco:time, ) give sending-efficient algorithms and claim that it is possible to extend the results in their papers so that each packet listens to the channel $O(\log N)$ times, with high probability, even without collision detection. However, these results deviate significantly from the ternary model (requiring control messages and, in the case of (DeMarcoSt17, ; marco:time, ), Good Samaritan packets).
We also note that full energy efficiency has drawn attention outside the domain of contention resolution. Many of these results address wireless networks where, in order to extend the network’s operational lifetime, low-power devices may spend significant time in an energy-efficient sleep state, waking up rarely to access the channel (e.g., (wang2019connectivity, ; deng2005scheduling, ; li2005energy, ; ye2006ultra, ; ullah2009study, ; merlin2010duty, )). Several algorithmic results address energy efficiency in the context of traditional distributed computing problems(e.g., (BarenboimM21, ; chatterjee2020sleeping, ; ghaffari2023distributed, ; dufoulon2023distributed, ; chang2018energy, ; king:sleeping, ; dani2021wake, )).
Dynamic packet arrivals. Much of the initial work on contention resolution addresses arrival times that are dictated by a stochastic process (MetcalfeBo76, ; Goldberg-notes-2000, ; hastad:analysis, ; goldberg2004bound, ; shoch1980measured, ; abramson1977throughput, ; kelly1987number, ; aldous:ultimate, ; GoodmanGrMaMa88, ; mosely1985class, ; RaghavanU98, ; RaghavanU95, ; GOLDBERG1999232, ). The intricacies of whether the contention-resolution algorithm knows local or global history, can adapt its actions over time, can listen on the channel, knows the number of packets in the system, etc. have been explored in depth (tsybakov1979ergodicity, ; ferguson1975control, ; fayolle1977stability, ; rosenkrantz1983instability, ; kelly:decentralized, ; paterson1995contention, ; DBLP:journals/jacm/GoldbergMPS00, ; tsybakov1978free-russian, ; tsybakov1978free, ; hayes:adaptive, ; massey1981collision, ; gallager, ; tsybakov1980random, ; verdu1986computation, ; tsybakov1987upper, ). A survey of much of the literature in this setting is provided by Chlebus (chlebus2001randomized, ).
In contrast to stochastic arrivals, packets may arrive at times scheduled by an adversary. The special case where all packets arrive simultaneously has been explored, and $O(n)$ makespan is possible (Gereb-GrausT92, ; GreenbergL85, ; FernandezAntaMoMu13, ; BenderFaHe05, ; BenderFiGi06, ). More generally, when packets are injected over time in a worst-case fashion, models from adversarial queuing theory (BorodinKlRa96, ; cruz-one, ; andrews2001universal, ) have been used to derive stability results (chlebus2009maximum, ; chlebus2007stability, ; anantharamu2010deterministic, ; aldawsari:adversarial, ; anantharamu2019packet, ; anantharamu2011medium, ; anantharamu2015broadcasting, ; hradovich2020contention, ; bienkowski:distributed, ; garncarek2018local, ; garncarek2019stable, ). Several results explore deterministic approaches for contention resolution (e.g., (deMarco:fast, ; greenberg1985lower, ; DBLP:conf/icdcs/MarcoKS19, ; DEMARCO20171, ; de2013contention, ; anantharamu2009adversarial, ; anantharamu2017adversarial, ; Kowalski05, ), although such approaches often have bounds that are asymptotically inferior to their corresponding randomized solutions.
Other computation on a shared channel. A closely-related problem to contention resolution is the wake-up problem, which addresses how long it takes until a single successful transmission (jurdzinski2015cost, ; jurdzinski2005probabilistic, ; jurdzinski2002probabilistic, ; jurdzinski2002probabilistic, ; farach2007initializing, ; farach2015initializing, ; Chlebus:2016:SWM:2882263.2882514, ; ChrobakGK07, ; ChlebusGKR05, ; DBLP:journals/dam/MarcoPS07, ; DBLP:conf/ictcs/MarcoPS05, ; willard:loglog, ; fineman:contention, ; fineman:contention2, ); this is in contrast to having all packets succeed.
More generally, several fundamental distributed computing challenges have been studied involving communication on a multiple access channel, and often even in the presence of adversarial jamming. These include broadcast (koo2, ; gilbert:near, ; gilbert:making, ; chen:broadcasting, ), leader election (gilbert:malicious, ; gilbert2009malicious, ; DBLP:journals/talg/AwerbuchRSSZ14, ), gossip (gilbert:interference, ; dolev:gossiping, ), node discovery (meier:speed, ), and simple point-to-point communication (king:conflict, ; DBLP:journals/dc/KingPSY18, ), packet scheduling (DBLP:conf/waoa/JurdzinskiKL14, ; DBLP:journals/tcs/AntaGKZ17, ).
7. Additional Discussion
We have provided a simple contention-resolution algorithm that achieves constant throughput with full energy efficiency (i.e., low sending and listening complexity), despite a powerful jamming adversary. This resolves in the affirmative two open questions about whether full-energy efficiency is possible at all in the popular ternary-feedback model, and whether it remains possible in the presence of jamming.
There are some natural directions for future work. The ability to prioritize network traffic is increasingly important to modern systems. For example, some real-time applications must guarantee message delivery within a bounded amount of time (WirelessHART, ; RoweMaRa06, ); additionally, assigning levels of priority to packets enhances fairness by preventing starvation (4411881, ; 4509751, ; 1286161, ). Agrawal et al. (DBLP:conf/spaa/AgrawalBFGY20, ) examine contention resolution when each packet has a deadline by which it must be successfully sent. However, their result is not listening efficient and handles only a form of random jamming. It may be interesting to explore whether jamming by a stronger adversary can be tolerated in a fully energy-efficient manner, where packets may be late, but only as a (slow-growing) function of the amount of jamming.
Another open problem is achieving similar guarantees in a multi-hop setting. Here, a major challenge is that packets listening in the same slot may receive different feedback. This aspect corresponds to the properties of signal propagation; for example, network participants may perceive different channel states depending on their distance from a sender. Richa et al. (richa:jamming2, ; richa:competitive-j, ) derive results for contention resolution in a multi-hop setting with jamming; however, their approach is not fully energy-efficient. The methods we employ here to achieve full energy efficiency would need revision in order to cope with the issue of packets receiving different channel feedback in the same slot.
Finally, it may be interesting to consider variable transmission strengths. For example, in certain cases where device hardware permits, jamming may be overcome by transmitting with sufficient energy. This would lead to a more complex cost model, where sending at a higher energy imposes a higher cost on the sender.
Appendix B
In this appendix, we include proofs for our preliminary results in Section 5.1, along with those for the observations in Section 5.2.
Proofs for Section 5.1
See 1
Proof.
To obtain the lower bound, note that the probability that some packet succeeds is at least:
$\sum_{\mbox{\tiny all packets $u$}}\left(p_{u}(t)\prod_{\mbox{\tiny all packets $v$}}(1-p_{v}(t))\right)\geq\frac{C(t)}{e^{2C(t)}}$
where the inequality follows from the left-hand side of Fact 1(a) and that $p_{v}\leq 1/2$ for all $v$, given that each packet always has a window size at least two.
To obtain the upper bound, we have:
$\sum_{\mbox{\tiny all packets $u$}}\left(p_{u}(t)\prod_{\mbox{\tiny all packets $v\not=u$}}(1-p_{v}(t))\right)$
$=$ $\sum_{\mbox{\tiny all packets $u$}}\left(p_{j}(t)\frac{(1-p_{u}(t))}{(1-p_{u}(t))}\prod_{\mbox{\tiny all packets $v\not=u$}}(1-p_{v}(t))\right)$
$\leq$ $e^{-C(t)}\left(\sum_{\mbox{\tiny all packets $u$}}\frac{p_{u}(t)}{1-p_{u}(t)}\right)$
$\leq$ $e^{-C(t)}2C(t),$
where the second-last line follows from applying the right-hand side Lemma 1, and the last line follows since $p_{v}\leq 1/2$ for all $v$. ∎
See 2
Proof.
Slot $t$ is empty when no packet sends in that slot, and this event occurs with probability:
$\prod_{\mbox{\tiny all packets u}}(1-1/w_{u}(t))$ $\geq e^{-2\sum_{\mbox{\tiny all packets u}}1/w_{u}(t)}$
$=e^{-2C(t)}.$
where the first inequality holds by Fact 1(a) and because window sizes are always at least $2$. The upper bound is derived as follows:
$\prod_{\mbox{\tiny all packets u}}(1-1/w_{u}(t))$ $\leq e^{-\sum_{\mbox{\tiny all packets u}}1/w_{u}(t)}$
$=e^{-C(t)}.$
where the first inequality holds by Fact 1(d). ∎
See 3
Proof.
The probability that slot $t$ contains noise is:
$1-p_{\rm suc}(t)-p_{\rm emp}(t)$ $\geq 1-\frac{2C(t)}{e^{C(t)}}-\frac{1}{e^{C(t)}}.$
Proofs for Section 5.2
See 7
Proof.
$H(t)$ is fixed (i.e., not a random variable) and each packet $u$ listens with probability $\frac{c\ln^{3}(w_{u}(t))}{w_{u}(t)}$. If $u$ listens to slot $t$, then $w_{u}(t+1)\leftarrow w_{u}(t)\left(1+\frac{1}{c\ln(w_{u}(t))}\right)$. Thus, we have:
$\mathbb{E}[H(t)-H(t+1)]$ $=\sum\frac{c\ln^{3}(w_{u}(t))}{w_{u}(t)}\left(\frac{1}{c\ln(w_{u}(t))}-\frac{1}{c\ln(w_{u}(t+1))}\right)$
$=\sum\frac{c\ln^{3}(w_{u}(t))}{w_{u}(t)}\left(\frac{1}{c\ln(w_{u}(t))}-\frac{1}{c\ln\left(w_{u}(t)\left(1+\frac{1}{c\ln(w_{u}(t))}\right)\right)}\right)$
$\geq\sum\frac{c\ln^{3}(w_{u}(t))}{w_{u}(t)}\left(\frac{1}{c\ln(w_{u}(t))}-\frac{1}{c\ln(w_{u}(t))+\frac{1}{\ln(w_{u}(t))}}\right)$
$=\sum\frac{c\ln^{3}(w_{u}(t))}{w_{u}(t)}\left(\frac{1}{c^{2}\ln^{3}(w_{u}(t))+c\ln(w_{u}(t))}\right)$
$\geq\sum_{u}\frac{c\ln^{3}(w_{u}(t))}{w_{u}(t)}\left(\frac{1}{2c^{2}\ln^{3}(w_{u}(t))}\right)$
$=\frac{C(t)}{2c}.$
The third line follows from Fact 1(b) using $x=\frac{1}{c\ln(w_{u}(t))}$ and noting that $c\ln\left(1+\frac{1}{c\ln(w_{u}(t))}\right)\leq\frac{1}{\ln(w_{u}(t))}$. By linearity of expectation, we have $\mathbb{E}[H(t+1)]\leq H(t)-\frac{C(t)}{2c}$, as claimed. ∎
See 8
Proof.
$H(t)$ is fixed and each packet $u$ listens with probability $c\ln^{3}(w_{u}(t))/w_{u}(t)$. If $u$ hears slot $t$, then $w_{u}(t+1)\leftarrow w_{u}(t)/\big{(}1+\frac{1}{c\ln(w_{u}(t))}\big{)}$. Thus, we have:
$\mathbb{E}[H(t+1)-H(t)]$ $=\sum\frac{c\ln^{3}(w_{u}(t))}{w_{u}(t)}\left(\frac{1}{c\ln(w_{u}(t+1))}-\frac{1}{c\ln(w_{u}(t))}\right)$
$=\sum\frac{c\ln^{3}(w_{u}(t))}{w_{u}(t)}\left(\frac{1}{c\ln\left(w_{u}(t)/\left(1+\frac{1}{c\ln w_{u}(t)}\right)\right)}-\frac{1}{c\ln w_{u}(t)}\right)$
$=\sum\frac{c\ln^{3}(w_{u}(t))}{w_{u}(t)}\left(\frac{1}{c\ln(w_{u}(t))-c\ln\left(1+\frac{1}{c\ln w_{u}(t)}\right)}-\frac{1}{c\ln(w_{u}(t))}\right)$
$\leq\sum\frac{c\ln^{3}(w_{u}(t))}{w_{u}(t)}\left(\frac{1}{c\ln(w_{u}(t))-\frac{1}{\ln(w_{u}(t))}}-\frac{1}{c\ln(w_{u}(t))}\right)$
$=\sum\frac{c\ln^{3}(w_{u}(t))}{w_{u}(t)}\left(\frac{1}{c^{2}\ln^{3}(w_{u}(t))-c\ln(w_{u}(t))}\right)$
$\leq\sum_{u}\frac{c\ln^{3}(w_{u}(t))}{w_{u}(t)}\left(\frac{2}{c^{2}\ln^{3}(w_{u}(t))}\right)$
$=\frac{2C(t)}{c}.$
The fourth line follows from Fact 1(b) using $x=\frac{1}{c\ln(w_{u}(t))}$. By linearity of expectation, we have $\mathbb{E}[H(t+1)]\leq H(t)+\frac{2C(t)}{c}$, as claimed. ∎