### v5

Abstract
Contention resolution addresses the problem of coordinating access to a shared communication channel. Time is discretized into synchronized slots, and a packet transmission can be made in any slot. A packet is successfully sent if no other packet is also transmitted during that slot. If two or more packets are sent in the same slot, then these packets collide and fail. Listening on the channel during a slot provides ternary feedback, indicating whether that slot had (0) silence, (1) a successful transmission, or (2+) noise. No other feedback or exchange of information is available to packets. Packets are (adversarially) injected into the system over time. A packet departs the system once it is successfully sent. The goal is to send all packets while optimizing throughput, which is roughly the fraction of successful slots.
Most prior contention resolution algorithms with constant throughput require a short feedback loop, in the sense that a packet’s sending probability in slot $t+1$ is fully determined by its internal state at slot $t$ and the channel feedback at slot $t$. This paper answers the question of whether these short feedback loops are necessary; that is, how often must listening and updating occur in order to achieve constant throughput? We can restate this question in terms of energy efficiency: given that both listening and sending consume significant energy, is it possible to have a contention-resolution algorithm with ternary feedback that is efficient for both operations?
A shared channel can also suffer random or adversarial noise, which causes any listener to hear noise, even when no packets are actually sent. Such noise arises due to hardware/software failures or malicious interference (all modeled as “jamming”), which can have a ruinous effect on the throughput and energy efficiency. How does noise affect our goal of long feedback loops/energy efficiency?
Tying these questions together, we ask: what does a contention-resolution algorithm have to sacrifice to reduce channel accesses? Must we give up on constant throughput? What about robustness to noise? Here, we show that we need not concede anything by presenting an algorithm with the following guarantees. Suppose there are $N$ packets arriving over time and $\mathcal{J}$ jammed slots, where the input is determined by an adaptive adversary. With high probability in $N+\mathcal{J}$, our algorithm guarantees $\Theta(1)$ throughput and $\operatorname{polylog}(N+\mathcal{J})$ channel accesses (sends or listens) per packet. We also have analogous guarantees when the input stream is infinite—we prove implicit throughput bounds of $\Omega(1)$ for all time slots $t$, and this translates to $\Theta(1)$ guaranteed throughput for any slot $t$ where the implicit throughput is sufficiently small in $\Theta(1)$. As a special case, these throughput results give rise to adversarial-queuing theory guarantees.
1 Introduction
Since the 1970s, randomized backoff protocols such as binary exponential backoff [MetcalfeBo76], have been used for managing contention on a shared communication channel. Originally used in the ALOHA system [ALOHAnet] and Ethernet [MetcalfeBo76], randomized backoff plays an important role in a wide range of applications, including WiFi [802.11-standard], wireless sensor networks [khanafer2013survey], transactional memory [scherer2005advanced, HerlihyMo93], and congestion control [wierman2003unified]. The salient feature of the communication channel is that it supports only one message transmission at a time: if more than one message is sent simultaneously, there is a collision resulting in indecipherable noise [Goldberg-notes-2000, kurose:computer, xiao:performance, DBLP:journals/jacm/GoldbergMPS00, MetcalfeBo76].
This contention-resolution problem is formalized as follows. There are $N$ packets arriving over time, and each packet needs to succeed (be the only packet sending) on the channel. [^1] For ease of exposition, we slightly abuse terminology and have the packets themselves taking action (e.g., sending themselves on the channel, listening on the channel), rather than introducing “agents”/“devices”/“senders”, where each one arrives on the scene with a packet to transmit. Time is divided into synchronized slots, each of which is sized to fit a single packet. To succeed, the packet requires exclusive access to the channel that is, the packet must be the only one transmitted during that slot. Otherwise, if two or more packets are transmitted in the same slot, the result is a collision, where none of the transmitted packets succeed. A packet departs the system once it succeeds. There is no a priori coordinator or central authority; packet transmissions are scheduled in a distributed manner. The objective is to have all packets succeed while optimizing the throughput of the channel, which is roughly the fraction of successful slots.
A popular contention-resolution protocol is binary exponential backoff [MetcalfeBo76] (see [kelly:decentralized, kelly1987number, rosenkrantz:some, aldous:ultimate, GoodmanGrMaMa88, GoodmanGrMaMa88, Al-Ammal2000, al2001improved, song:stability]). Informally, under binary exponential backoff, a packet that has been in the system for $t$ slots is sent with probability $\Theta(1/t)$.
Feedback loops: short versus long (versus no feedback).
An elegant, but ultimately problematic, feature of classical binary exponential backoff is that it is oblivious—packets remaining in the system do not use channel feedback to adjust their behavior: a packet with age $t$ sends with probability $\Theta(1/t)$ until it succeeds, regardless of channel history. The unfortunate result is that with adversarial packet arrivals, binary exponential backoff supports only a subconstant throughput—specifically, $O(1/\ln{N})$; in fact, even for the batch case where all $N$ packets arrive at the same time, the throughput of binary exponential backoff is only $O(1/\ln N)$ [BenderFaHe05].
In contrast, contention-resolution protocols can exploit frequent channel feedback to achieve $\Theta(1)$ throughput under adversarial arrivals [BenderKoKu20, BenderFiGi19, BenderKoPe16, ChangJiPe19, awerbuch:jamming]. These protocols are not oblivious—packets listen on the channel and adjust their sending probabilities up or down based on this feedback.
The primary model for channel feedback is the ternary-feedback model [BenderFiGi19, DBLP:conf/spaa/AgrawalBFGY20, gilbert:near, DBLP:journals/dc/KingPSY18, chlebus2009maximum, chlebus2007stability, ChlebusKoRo06, fineman:contention2, anantharamu2019packet, greenberg1985lower]. In this model, a packet can listen on the channel in each slot and learn whether that slot is (0) empty, if no packets send, (1) successful, if exactly one packet sends, or (2+) noisy, if two or more packets send. Based on this feedback, the packet can decide when to attempt to send next. Once a packet succeeds, it immediately departs the system.
Most constant-throughput algorithms with ternary-feedback (e.g., [BenderKoKu20, BenderFiGi19, BenderKoPe16, ChangJiPe19, DeMarcoSt17, awerbuch:jamming, richa:jamming2, richa:jamming4, richa:jamming3, ogierman:competitive]) listen on the channel in every slot (or every constant number of slots). That is, these algorithms have a short feedback loop: a packet’s sending probability in slot $t+1$ is fully determined by its internal state at slot $t$ and the channel feedback at slot $t$. For example, the algorithm by Chang, Jin, and Pettie [ChangJiPe19], listens in every slot $t$ and multiplicatively updates the sending probability in slot $t+1$ based on whether it heard silence, a successful transmission, or a collision in slot $t$.
A key question is as follows: how frequently does the packet need to listen on the channel and update its behavior in order to achieve constant throughput? In every slot? In a vanishingly small fraction of slots? How short a feedback loop is necessary for good throughput?
As an analogy, one cannot navigate a ship without a control feedback loop: monitoring the surroundings and correcting course. One option is to constantly monitor and continuously update the heading to avoid obstacles. But the analogous question is whether one can still safely navigate with only a vanishingly small amount of course correction. In contention resolution, the monitoring is via the channel sensing and the course being corrected is the transmission probabilities.
Robustness to noise.
Finally, in addition to the factors discussed above, much of the recent work on contention resolution (see [anantharamu2019packet, Chlebus:2016:SWM:2882263.2882514, DBLP:conf/spaa/AgrawalBFGY20, ChenJiZh21, jiang:robust, awerbuch:jamming, richa:jamming2, richa:jamming4, richa:jamming3, ogierman:competitive, BenderFiGi19, ChangJiPe19]) has sought to address an additional factor, which is that the real world is often noisy. Sometimes interference prevents transmissions in a slot and listeners hear noise even if nobody actually sends. Noisy channels arise due to hardware/software failures, co-located devices, or malicious jamming [chiwewe2015using, liu2017interference, pelechrinis:denial, mpitziopoulos:survey, xu:feasibility]. Regardless of the source of the noise in a slot, we may think of these slots as being “jammed” by an adversary. Jamming has evolved from a mostly-theoretical risk into a credible threat to systems over the past decade, with several publicized examples [nicholl, FCC:humphreys, marriot, priest].
After a long line of work, there are now many contention-resolution protocols that achieve constant throughput in the presence of noise (e.g. [awerbuch:jamming, richa:jamming2, richa:jamming4, richa:jamming3, ogierman:competitive, BenderFiGi19, ChangJiPe19]); however, again, these protocols listen to the channel in every slot and update their sending probabilities accordingly. Our goal is to eliminate short feedback loops not just in the classical model, but also with jamming.
Minimizing channel accesses = energy efficiency.
Up until now, we have discussed whether one can minimize listening and thus avoid providing immediate feedback for a contention-resolution algorithm. Another way of viewing this problem is through the lens of energy efficiency. Each channel access—whether for sending or listening—consumes energy. Most work on contention resolution is sending efficient, but is not listening efficient (e.g.,  [awerbuch:jamming, richa:jamming2, richa:jamming3, richa:jamming4, BenderFiGi19, BenderKoKu20, DeMarcoSt17, jiang:robust, de2022contention]). That is, most protocols optimize how frequently a packet sends, but allow a packet to listen in every slot “for free”.
In fact, both sending and listening are expensive operations (e.g., [polastre:telos, feeney2001investigating, yang2017narrowband, lauridsen2018empirical]), which should be minimized if devices are energy-constrained (e.g., battery-powered devices). Minimizing energy usage by having devices sleep as much as possible has been a long-standing and popular strategy to maximize the network lifetime (for example, the development of duty-cycle protocols [merlin2010duty, ye2006ultra, li2005energy]). Why optimize listening in contention resolution protocols if packets must receive messages (for other network applications)? Informally, there are two types of listening: listening to receive messages, and listening to execute a contention resolution protocol (which is what enables sending messages). To receive messages, packets do not need to listen in every time slot; there exist methods for optimizing this first type of listening. Although this is beyond the scope of this paper, as one example, in many wireless settings, a central base station can monitor the channel constantly and facilitate message exchange [feng2012survey].
We call a protocol fully energy-efficient if it is both sending efficient and listening efficient. By definition, such a protocol cannot have short feedback loops, since it can access the channel only rarely.
Past work: Minimizing listening by allowing for explicit synchronization.
Currently, the only known path to full energy efficiency is via explicit synchronization. This means that the model is extended so that packets can send synchronization messages to each other whenever they broadcast [BenderKoPe16, DeMarcoSt17, marco:time]. In [BenderKoPe16], these synchronization messages have size $\Theta(\ln N)$ bits each, which means that an arbitrary polynomial amount of communication can be expressed in a slot. In [DeMarcoSt17, marco:time], the synchronization messages are smaller ($O(1)$ bits), but some packets are permitted to stick around as “Good Samaritans” in order to serve as long-term coordinators (that send many messages via many broadcasts). We highlight that the model used in [DeMarcoSt17, marco:time] is directly comparable to our model. The language used in those works frames the problem as devices attempting to successfully send packets on a shared channel. In that language, the devices are allowed to remain in the system issuing additional coordination messages after successfully sending their packet.
Using $\Theta(\ln N)$-bit synchronization messages, Bender, Kopelowitz, Pettie, and Young [BenderKoPe16] give an algorithm with $\Theta(1)$ throughput and expected $O(\ln(\ln^{*}N))$ channel accesses per packet. Using Good-Samaritan packets, De Marco and Stachowiak [DeMarcoSt17] and De Marco, Kowalski, and Stachowiak [marco:time] provide a constant-throughput algorithm that is sending efficient ($O(\ln N)$ transmissions per packet) and conjecture that their techniques can be extended to achieve fully energy efficiency ($O(\operatorname{polylog}N)$ channel accesses per packet), with high probability and even without collision detection.
In all of these cases, even with the help of explicit synchronization, it remains open whether one can achieve such results in the presence of adversarial noise. Indeed, noise has the potential to be extra-problematic for energy-efficient algorithms since these algorithms listen to the channel less frequently and can potentially be thrown off by a small amount of well-placed noise.
This paper.
We show that it is indeed possible to achieve constant throughput with full energy efficiency while being robust to adversarial noise. Moreover, these results hold in the standard ternary-feedback model, without requiring the addition of any sort of explicit synchronization.
Our algorithm belongs to a natural family of multiplicative-weight-update algorithms (e.g., [ChangJiPe19, awerbuch:jamming, richa:jamming2, richa:jamming3, ogierman:competitive, richa:jamming4]). When a packet hears silence, it multiplicatively increases both its listening and sending probabilities. Conversely, when a packet hears noise, it multiplicatively decreases these probabilities. There are no control messages, no Good Samaritan packets, and no leaders elected.
What makes our algorithm/analysis interesting is that we are able to support a multiplicative-weight-update framework while having each packet ‘cover its eyes’ almost all of the time. This is in stark contrast to prior constant-throughput algorithms, which adjust sending probabilities in every slot. Because each packet listens to so few slots, the different packets that are in the system at the same time may end up with very different perspectives on the world from each other. In order to analyze the ‘herd behavior’ of the packets in this potentially chaotic setting, substantially new techniques end up being required. These techniques are also what allow us to handle the additional chaos that adversarial jamming adds to the system.
1.1 Model
A finite or infinite stream of indistinguishable packets arrives over time; the number of arrivals is unknown to the algorithm. Each packet must be sent on the multiple-access channel. Time is divided into synchronized slots, each of which is sufficiently large to send a single packet. An adversary (specified below) controls how many packets are injected into the system in each slot. When a packet succeeds, it departs the system. There is no universal numbering scheme for the slots; that is, there is no global clock from which a packet could infer the system lifetime or slot parity. Additionally, the packets do not receive any additional information about how many packets have arrived or will arrive. Instead, packets only receive information through the ternary feedback model.
We now describe the ternary feedback model. In each time slot, each packet in the system can take one of three actions: (i) sleep, (ii) send, or (iii) listen to the channel. Packets that take actions (ii) or (iii) are said to access the channel. If no packets choose to send during a slot, then that (non-jammed) slot is empty/silent. A packet that listens during a slot (action iii) learns whether the slot was (0) empty, (1) successful, or (2+) noisy. A packet that sleeps during a slot (action i) learns nothing about the state of the slot. A packet that sends (action ii), either succeeds and leaves the system, or collides and remains in the system. [^2] For ease of presentation, in our algorithms, we say that a packet can listen and send simultaneously, but any packet that is sending actually does not need to listen to determine the state of the channel. If the packet is still in the system after sending in slot $t$, then slot $t$ was noisy.
We now add jamming to the picture; an adversary determines which slots are jammed. To jam a particular slot, the adversary broadcasts noise into that slot. All jammed slots are thus full and noisy. A packet that listens in a jammed slot hears that the slot was noisy, but does not know whether that noise came from jamming or was merely a collision between two or more packets. A packet that sends during a jammed slot collides and thus remains in the system.
An adversary determines, for each slot $t$, how many packets to inject in slot $t$ and whether to jam in that slot. This paper considers an adaptive adversary, which bases its decision on the entire state of the system so far, i.e., up to the end of slot $t-1$, but not the outcomes of future coin tosses. Thus, if at slot $t$, a packet $p$ decides whether to send based on a coin flip, the adaptive adversary does not get to see that coin flip until after slot $t$.
A basic metric: (overall) throughput.
The main objective of contention resolution is to optimize throughput, defined next. A slot is active if at least one packet is in the system during that slot; inactive slots can be ignored in our analysis. Similarly, a packet is active if has been injected into the system and has not yet departed. Without loss of generality, assume throughout that the first slot is active. Without jamming, the throughput at time $t$ is defined as $T_{t}/S_{t}$, where $T_{t}$ is the number of successful transmissions during slots $1,2,\ldots,t$, and $S_{t}$ is the number of active slots in the same time interval. [^3] By assumption that the first slot is active, we have $S_{t}\geq 1$ and hence throughput is always well defined. On a finite input, the (overall) throughput is defined with respect to the final active slot $t$; at this point, the overall throughput is $N/S$, where $N=T_{t}$ is the total number of packets and $S=S_{t}$ is the total number of active slots. Overall throughput is not well-defined on an infinite execution.
A deficiency of the throughput metric (when defined naïvely) is that even if an algorithm guarantees $\Theta(1)$ overall throughput, it is not possible to achieve $\Theta(1)$ throughput uniformly across time. For example, if there is a burst of $N$ packets at time 0, and $N$ is unknown to the algorithm, then the throughput will be 0 for a superconstant number of slots (e.g., see [willard:loglog, ChangKoPe19]), and this is provably unavoidable regardless of the backoff strategy being used. Thus (overall) throughput is only meaningful at the end of the execution, or at points in time where there are no packets in the system.
A stronger metric: implicit throughput [BenderKoKu20].
In order to support a meaningful notion of throughput, even at intermediate points in time, Bender, Kopelowitz, Kuszmaul, and Pettie [BenderKoKu20] propose a refined definition that they term “implicit throughput” [BenderKoKu20]. The implicit throughput at time $t$ is defined as $N_{t}/S_{t}$, where $N_{t}$ is the total number of packets that arrive at or before time $t$ and $S_{t}$ is the total number of active slots so far. [^4] Notice that while $N_{t}$ depends on the adversary, the number of active slots depends on the algorithm, whose goal it is to make slots inactive by completing packets.
One perspective on implicit throughput is that it is an analytical tool. Indeed, whenever we reach a point in time where overall throughput is meaningful (i.e., there are no packets left in the system), the two metrics become provably equal. This includes both at the end of any finite execution or during quiet periods of infinite executions. Another perspective on implicit throughput is that it is a stronger metric that offers a meaningful guarantee even at intermediate points in time: what constant implicit throughput means is that the number of active slots used so far should never be asymptotically larger than the number of packets that have arrived.
Since we ignore slots where no packets are in the system, the implicit throughput of our algorithm (see Section 3) is guaranteed (w.h.p.) to be $\Omega(1)$. Moreover, for every slot $t$ where the implicit throughput up to that slot is sufficiently small in $\Theta(1)$, the standard notion of throughput is also $\Theta(1)$.
Throughout the rest of the paper, we focus exclusively on implicit throughput. We should emphasize, however, that this only makes our results stronger—the results also imply the standard constant-throughput guarantees that one would normally strive for.
Extending to adversarial jamming.
We next extend the definitions of throughput and implicit throughput for the case of jamming following [BenderFiGi19]. An algorithm wastes a slot if that slot has silence or a collision, and throughput measures the fraction of slots that the algorithm could have used but instead wasted. Let $\mathcal{J}_{t}$ denote the number of jammed slots through slot $t$. Then, the throughput of an execution ending at time $t$ is defined to be $(T_{t}+\mathcal{J}_{t})/S_{t}$, and the implicit throughput at slot $t$ is defined to be $(N_{t}+\mathcal{J}_{t})/S_{t}$.
Some useful properties of implicit throughput, and applications to adversarial queuing theory.
We conclude the section by summarizing several useful properties of implicit throughput.
[…]
Consider any inactive slot $t$, i.e., where there are no active packets in the system. Then the implicit throughput and throughput are the same at slot $t$.
[…]
Let $\delta$ be any lower bound on the implicit throughput of an algorithm; that is, suppose the algorithm achieves implicit throughput of at least $\delta$ at all times. Let $N_{t}$ and $S_{t}$ denote the total number of packet arrivals and active slots, respectively, at or before time $t$. Then $S_{t}\leq N_{t}/\delta$. Consequently:
  • Overall throughput. Suppose that there are a total of $N\geq 1$ packet arrivals. Then the total number of active slots is at most $N/\delta$, and hence the overall throughput is at least $\delta$.
  • Backlog reprieve. If $N_{t}<\delta t$, then there exists an inactive slot $t^{\prime}\leq t$. Thus, all packets that arrived before $t^{\prime}$ have completed before $t^{\prime}$, and hence the throughput at time $t^{\prime}$ is at least $\delta$.
Finally, we observe that there are several natural settings in which implicit-throughput guarantees directly imply strong guarantees on packet backlog, even for infinite input sequences. Suppose, in particular, that packets arrive according to the adversarial queuing theory arrival model (described below), which parameterizes the “burstiness” of packet-arrival in infinite streams. In adversarial queuing theory [BenderFaHe05, ChlebusKoRo12, ChlebusKoRo06, BorodinKlRaSu01], the adversary is restricted from injecting too many packets/jammed slots over a set of consecutive slots of length $S$, where $S$ is a parameter of the model that we refer to as the granularity. For granularity $S$, the number of packet arrivals plus the number of jammed slots is limited to $\lambda S$, where the arrival rate $\lambda$ is a constant less than one. On the other hand, how the packet arrivals are distributed within each $S$-sized window is adversarial: no restrictions are placed on how the (at most) $\lambda S$ packets and jammed slots are distributed. By showing that the implicit throughput of all active slots is $\Omega(1)$, we obtain as a corollary a strong bound of $O(S)$ on the number of packets backlogged in the system at all times, as long as $\lambda$ is a sufficiently small constant.
1.2 Main results
We present an algorithm that with high-probability guarantees $\Omega(1)$ implicit throughput in all active slots and full energy efficiency. Thus, we resolve two open questions in contention resolution: First, we show that full-energy efficiency is feasible, even with only ternary feedback. Second, we show that these guarantees are achievable, even in the presence of adversarial jamming.
Our algorithm relies on the natural multiplicative-weight approach to backoff—with a careful choice of probabilities and update rules. In contrast to some previous approaches, we do not rely on packet batching (to turn the online problem into a series of batch problems), leader election, busy tones, population estimation, dividing the channel into simulated subchannels (e.g., odd and even slots) for packet coordination, or other approaches seen in many modern algorithms; alas, these seem problematic in our setting. Given the simplicity of our algorithm, the technical innovation lies in choosing the update parameters, analyzing the underlying combinatorial process, and proving that it is fast, robust, and fully energy-efficient.
Our algorithm Low-Sensing Backoff guarantees the main theorems below. [^5] An event occurs with high probability (w.h.p.) in $x$ if for any fixed constant $c\geq 1$, the probability of the event is at least $1-x^{-c}$.
Theorem 1.1 (Implicit throughput, infinite packet streams).
At the $t$-th active slot, the implicit throughput is $\Omega(1)$ w.h.p. in $t$.
Corollary 1.2 (Throughput, finite packet streams).
Consider an input stream of $N$ packets with $\mathcal{J}$ jammed slots. The throughput for the execution is $\Theta(1)$ w.h.p. in $N+\mathcal{J}$.
Corollary 1.3 (Bounded backlog for adversarial-queuing arrivals).
Consider adversarial-queuing-theory arrivals with a sufficiently small constant arrival-rate $\lambda$ and granularity $S$; i.e., in any interval of length $S$, the total number of packet arrivals and jammed slots is at most $\lambda S$. Then, for any given slot $t$, the number of packets currently in the system is at most $O(S)$ w.h.p. in $S$.
Theorem 1.4 (Energy, finite executions).
Consider a finite execution with $N$ total packet arrivals and $\mathcal{J}$ total jammed slots against an adaptive adversary. Any given packet accesses the channel $O(\operatorname{polylog}$ $(N+\mathcal{J}))$ times w.h.p. in $N+\mathcal{J}$.
Theorem 1.5 (Energy, adversarial queuing).
Consider a (finite or infinite) packet stream with adversarial-queuing arrivals with granularity $S$ and where the arrival rate is a sufficiently small constant. Then any given packet accesses the channel $O(\operatorname{polylog}(S))$ times w.h.p. in $S$ against an adaptive adversary.
Theorem 1.6 (Energy, infinite executions).
Consider an infinite packet stream, and let $N_{t}$ and $\mathcal{J}_{t}$ denote the number of arrivals and jammed slots, respectively, up until time $t$ against an adaptive adversary. Then any given packet accesses the channel $O(\operatorname{polylog}(N_{t}+\mathcal{J}_{t}))$ times before time $t$, w.h.p. in $N_{t}+\mathcal{J}_{t}$.
The above statements are simplified slightly for ease of presentation, while the formal statements are presented later in Section LABEL:sec:analysis. Specifically, Theorem 1.1 is proved in Section LABEL:sec:bettinggame where it appears as Corollary LABEL:cor:implicit-throughput, and Corollary 1.3 corresponds to Corollary LABEL:cor:backlog. The remaining theorems are proved in Section LABEL:sec:energy. In particular, Theorem 1.4 corresponds to Theorem LABEL:thm:finite_adaptive_accesses, and Theorem 1.5 corresponds to Theorem LABEL:thm:queuing_adversarial_accesses. Finally, Theorem 1.6 is included in Theorem LABEL:thm:infinite_accesses.
1.3 Extensions to Reactive Adversary
A
Definition 1.
reactive adversary [richa:jamming3, wilhelm2011short, DBLP:journals/dc/KingPSY18] has an instantaneous reaction time; that is, this adversary listens to the channel and can decide whether to jam and/or inject new packet(s) in slot $t$ based on what it hears in slot $t$ itself. In contrast, the standard adaptive adversary would not know whether any packet chooses to send in slot $t$ until slot $t+1$. This allows a reactive adversary to cheaply prevent any particular packet $p$ from succeeding by jamming only those slots where $p$ makes transmission attempts. Thus, against a reactive adversary, the total number of channel accesses required is at least linear in the amount of jamming. For exponential backoff, the situation is more dire: for any $T$ a reactive adversary can also drive the throughput down to $O(1/T)$ by jamming a single packet a mere $\Theta(\ln T)$ times.
Note that reactivity and adaptivity are somewhat orthogonal. Reactivity addresses how quickly the adversary can react to the detectable channel state—importantly, only sending is revealed, since it is detectable. In contrast, an adaptive adversary knows all of the internal state and random choices of packets up to the previous slot, and in particular this adversary also knows if packets choose to listen.
In addition to our main results on purely adaptive adversaries, we also address an adversary that is both adaptive and reactive. It turns out that the reactive adversary does not impact our implicit-throughput bounds for our algorithm (our analysis applies whether or not the adversary can see the channel activity at the current time). Reactivity thus only impacts the number of channel accesses. Roughly speaking, the theorem states that the reactive adversary has nontrivial impact on the worst-case number of channel accesses and thus energy, which is to be expected, but it does not have significant impact on the average.
Theorem 1.7 (Energy, reactive adversary).
The following apply to a reactive and adaptive adversary.
  1. 1.
    Finite streams. Consider a finite execution with $N$ total packet arrivals and $\mathcal{J}$ total jammed slots. Any given packet accesses the channel $O((\mathcal{J}+1)\operatorname{polylog}(N))$ times w.h.p. in $N+\mathcal{J}$. Moreover, the average number of channel accesses is only $O((\mathcal{J}/N+1)\operatorname{polylog}(N+\mathcal{J}))$ times w.h.p. in $N+\mathcal{J}$.
  2. 2.
    Adversarial queuing. Consider a (finite or infinite) packet stream with adversarial-queuing arrivals with granularity $S$ and where the arrival rate is a sufficiently small constant. Then any given packet accesses the channel at most $O(S)$ times, w.h.p. in $S$. In addition, the average number of accesses per slot is $O(\operatorname{polylog}(S))$, w.h.p. in $S$.
  3. 3.
    Infinite executions. Consider an infinite packet stream, and let $N_{t}$ and $\mathcal{J}_{t}$ denote the number of arrivals and jammed slots, respectively, up until time $t$. Then any given packet accesses the channel $O((\mathcal{J}_{t}+1)\operatorname{polylog}(N_{t}+\mathcal{J}_{t}))$ times before time $t$, w.h.p. in $N_{t}+\mathcal{J}_{t}$. Moreover, the average number of channel accesses is $O((\mathcal{J}_{t}/N_{t}+1)\operatorname{polylog}(N_{t}+\mathcal{J}_{t}))$.
The items of Theorem 1.7 are each proved separately as Theorems LABEL:thm:finite_reactive_accesses, LABEL:thm:queuing_reactive_accesses, and LABEL:thm:infinite_accesses in Section LABEL:sec:energy.
2 Related Work
Contention resolution has been an active area of research for several decades. Here, we discuss related results from the literature. We highlight that a preliminary version of this work has appeared previously [bender2024fully].
Robustness.
Early work on jamming resistance by Awerbuch, Richa, and Scheideler [awerbuch:jamming] considers the case of $N$ devices, each of which always has a packet to ready to send; that is, rather than changing over time, the number of packets contending for channel access is always $N$. For any interval of at least $T$ consecutive slots, at most a $(1-\lambda)$-fraction may be jammed, where the constant $\lambda>0$ is unknown. In this setting, Awerbuch, Richa, and Scheideler[awerbuch:jamming] showed that constant throughput is possible with a polylogarithmic number of transmissions per packet. Several subsequent results have built on this result, addressing multi-hop [richa:jamming2, richa:competitive-j] and co-located networks [richa:jamming4], reactive jamming [richa:jamming3, richa:efficient-j], and aspects of signal propagation [ogierman:competitive]. However, this series of results relies on devices knowing rough bounds on $N$ and $T$; in particular, devices must set a parameter whose value is $O(1/(\ln\ln(N)+\ln(T)))$.
A similar restriction on jamming is explored by Anantharamu et al. [anantharamu2019packet, anantharamu2011medium], where $\rho t+b$ slots in any window of consecutive $t$ slots can be jammed, for $0<\rho\leq 1$, and a non-negative integer, $b$. Probabilistic jamming has also been considered by Chlebus et al. [Chlebus:2016:SWM:2882263.2882514] in a multi-channel setting, and by Agrawal et al. [DBLP:conf/spaa/AgrawalBFGY20] where packets have delivery deadlines.
When the amount of jamming can be unbounded, Bender et al. [BenderFiGi19, BenderFiGi16] showed expected constant throughput is possible over non-jammed slots using a $\operatorname{polylog}(N)$ number of sending attempts per packet. Chang, Jin, and Pettie [ChangJiPe19] give an elegant algorithm that achieves the same asymptotic result, but drastically improves the expected throughput, rivaling the $1/e$ result achieved by the well-known slotted ALOHA [binder:aloha], despite the more-challenging adversarial setting.
There is a growing body of work [ChenJiZh21, jiang:robust, DeMarcoSt17, BenderKoKu20] on how to design contention resolution protocols where packets listening on the channel during a given slot learn whether or not that slot was successful, but not whether the slot was empty or noisy; that is, packets lack the ability to detect collisions. As mentioned earlier, the results by De Marco and Stachowiak [DeMarcoSt17], De Marco, Kowalski, and Stachowiak [marco:time], and Bender et al. [BenderKoKu20] address this setting. Dealing with jamming in the absence of collision detection is challenging. Here, results by Chen, Jiang, and Zheng [ChenJiZh21] and Jiang and Zheng [jiang:robust] show how jamming can still be tolerated, although throughput is $\Theta(1/\ln N)$ and packets are allowed to listen on the channel for free.
There is a large body of work on applied security considerations with respect to jamming. The impact of these attacks has been evaluated under a variety of wireless settings and adversarial strategies [bayraktaroglu:performance, xu:feasibility, lin:link, aschenbruck:simulative], and several defenses have been proposed. For example, spread-spectrum techniques involve devices coordinating their communication over different portions of the spectrum in order to evade collisions [spread:foiling, navda:using, deejam:wood]. Another defense is to have devices map the areas impacted by jamming and then routing traffic around those areas [wood2003jam, han2018hierarchical]. Specialized antenna technology can be used to filter out a jamming signal, which is a process referred to as null steering [zheng2018design, mcmilin2015gps].
Energy efficiency.
As discussed in Section 1, there are a handful of results on contention resolution that achieve constant throughput and address full energy efficiency. To reiterate, Bender et al. [BenderKoPe18] achieve expected $O(\ln(\ln^{*}N))$ channel accesses per packet, which Chang et al. [ChangKoPe19] prove is tight. De Marco and Stachowiak [DeMarcoSt17] and De Marco, Kowalski, and Stachowiak [marco:time] give sending-efficient algorithms and conjecture that it is possible to extend the results in their papers so that each packet listens to the channel a polylogarithmic number of times (see Section 7 Discussion and open problems in [marco:time]). However, these results deviate significantly from the ternary model (requiring control messages and, in the case of [DeMarcoSt17, marco:time], Good Samaritan packets).
We also note that full energy efficiency has drawn attention outside the domain of contention resolution. Many of these results address wireless networks where, in order to extend the network’s operational lifetime, low-power devices may spend significant time in an energy-efficient sleep state, waking up rarely to access the channel (e.g., [wang2019connectivity, deng2005scheduling, li2005energy, ye2006ultra, ullah2009study, merlin2010duty]). Several algorithmic results address energy efficiency in the context of traditional distributed computing problems(e.g., [BarenboimM21, chatterjee2020sleeping, ghaffari2023distributed, dufoulon2023distributed, chang2018energy, king:sleeping, dani2021wake]).
Dynamic packet arrivals.
Much of the initial work on contention resolution addresses arrival times that are dictated by a stochastic process [MetcalfeBo76, Goldberg-notes-2000, hastad:analysis, goldberg2004bound, shoch1980measured, abramson1977throughput, kelly1987number, aldous:ultimate, GoodmanGrMaMa88, mosely1985class, RaghavanU98, RaghavanU95, GOLDBERG1999232]. The intricacies of whether the contention-resolution algorithm knows local or global history, can adapt its actions over time, can listen on the channel, knows the number of packets in the system, etc. have been explored in depth [tsybakov1979ergodicity, ferguson1975control, fayolle1977stability, rosenkrantz1983instability, kelly:decentralized, paterson1995contention, DBLP:journals/jacm/GoldbergMPS00, tsybakov1978free-russian, tsybakov1978free, hayes:adaptive, massey1981collision, gallager, tsybakov1980random, verdu1986computation, tsybakov1987upper]. A survey of much of the literature in this setting is provided by Chlebus [chlebus2001randomized].
In contrast to stochastic arrivals, packets may arrive at times scheduled by an adversary. The special case where all packets arrive simultaneously has been explored, and $O(n)$ makespan is possible [Gereb-GrausT92, GreenbergL85, FernandezAntaMoMu13, BenderFaHe05, BenderFiGi06]. More generally, when packets are injected over time in a worst-case fashion, models from adversarial queuing theory [BorodinKlRa96, cruz-one, andrews2001universal] have been used to derive stability results [chlebus2009maximum, chlebus2007stability, anantharamu2010deterministic, aldawsari:adversarial, anantharamu2019packet, anantharamu2011medium, anantharamu2015broadcasting, hradovich2020contention, bienkowski:distributed, garncarek2018local, garncarek2019stable]. Several results explore deterministic approaches for contention resolution (e.g., [deMarco:fast, greenberg1985lower, DBLP:conf/icdcs/MarcoKS19, DEMARCO20171, de2013contention, anantharamu2009adversarial, anantharamu2017adversarial, Kowalski05], although such approaches often have bounds that are asymptotically inferior to their corresponding randomized solutions.
Other computation on a shared channel.
A closely-related problem to contention resolution is the wake-up problem, which addresses how long it takes until a single successful transmission [jurdzinski2015cost, jurdzinski2005probabilistic, jurdzinski2002probabilistic, jurdzinski2002probabilistic, farach2007initializing, farach2015initializing, Chlebus:2016:SWM:2882263.2882514, ChrobakGK07, ChlebusGKR05, DBLP:journals/dam/MarcoPS07, DBLP:conf/ictcs/MarcoPS05, willard:loglog, fineman:contention, fineman:contention2]; this is in contrast to having all packets succeed.
More generally, several fundamental distributed computing challenges have been studied involving communication on a multiple access channel, and often even in the presence of adversarial jamming. These include broadcast [koo2, gilbert:near, gilbert:making, chen:broadcasting], leader election [gilbert:malicious, gilbert2009malicious, DBLP:journals/talg/AwerbuchRSSZ14], gossip [gilbert:interference, dolev:gossiping], node discovery [meier:speed], and simple point-to-point communication [king:conflict, DBLP:journals/dc/KingPSY18], packet scheduling [DBLP:conf/waoa/JurdzinskiKL14, DBLP:journals/tcs/AntaGKZ17].
3 Low-Sensing Backoff Algorithm
This section presents the Low-Sensing Backoff algorithm; see Figure 1. For ease of presentation, we describe our algorithm as listening whenever it sends. However, the packet need not actually do both; observe that any packet that is sending does not need to listen to determine the state of the channel, since if the packet is still in the system after sending in slot $t$, then slot $t$ was noisy.
Low-Sensing Backoff for packet $u$ Key Variables: $w_{u}(t)$: window size of $u$ in slot $t$.
If $u$ is injected at time slot $t$, then $w_{u}(t)=w_{\rm min}$.
$c$: a sufficiently large positive constant.
In every slot $t$, packet $u$ executes the following four steps with probability $\frac{c\ln^{3}\left(w_{u}(t)\right)}{w_{u}(t)}$: Listen Send with probability $\frac{1}{c\ln^{3}(w_{u}(t))}$ If $u$ heard a silent slot, then $w_{u}(t+1)\leftarrow\max{\left\{\mbox{\small$\displaystyle\frac{w_{u}(t)}{1+1/(c\ln(w_{u}(t)))}$},\,w_{\rm min}\right\}}$ If $u$ heard a noisy slot, then $w_{u}(t+1)\leftarrow w_{u}(t)\cdot\left(\mbox{\small$\displaystyle 1+\frac{1}{c\ln(w_{u}(t))}$}\right)$
Figure 1: Low-Sensing Backoff algorithm.
The probabilities for sending and listening in Low-Sensing Backoff are determined by a single parameter, which we call